### v1

# HarmonicEval: Multi-modal, Multi-task, Multi-criteria Automatic Evaluation Using a Vision Language Model

Abstract
Vision-language models (VLMs) have shown impressive abilities in text and image understanding. However, existing metrics for evaluating the text generated by VLMs focus exclusively on overall quality, leading to two limitations: 1) it is challenging to identify which aspects of the text need improvement from the overall score; 2) metrics may overlook specific evaluation criteria when predicting an overall score. To address these limitations, we propose HarmonicEval, a reference-free evaluation metric that aggregates criterion-wise scores to produce the overall score in a bottom-up manner. Furthermore, we construct the Multi-task Multi-criteria Human Evaluation (MMHE) dataset, which comprises 18,000 expert human judgments across four vision-language tasks. Our experiments demonstrate that HarmonicEval achieves higher correlations with human judgments than conventional metrics while providing numerical scores for each criterion.
1 Introduction
Automatic evaluation of text generated by vision-language models (VLMs) is essential for enhancing their robustness and reliability in tasks such as image captioning and visual question answering Hessel et al. (2021); Lee et al. (2024b). As VLMs advance rapidly, the need for more informative and accurate metrics aligned with human perception has grown. However, most existing metrics focus solely on measuring the overall quality of texts generated by VLMs, lacking two critical aspects expected in automatic evaluation: 1) numerical explainability and 2) comprehensiveness.
Figure 1: Necessity of multi-criteria evaluation. (a) Example of text quality evaluation in referring expression generation. HarmonicEval provides scores for each criterion. (b) Example in visual question answering where existing interpretable approach overlooks conciseness.
First, evaluation metrics that provide only overall scores Papineni et al. (2002); Zhang et al. (2020); Hessel et al. (2021) cannot identify specific deficiencies in aspects of the text, such as clarity or fluency, as illustrated in Figure 1(a). While recent metrics for vision-language tasks Chan et al. (2023); Lee et al. (2024b) offer free-form textual explanations to complement scores, these explanations lack the granularity and consistency that come from explicit numerical breakdowns. As a result, providing a numerical explanation that links specific aspect qualities to overall scores remains a challenge in the automatic evaluation of vision-language tasks.
Figure 2: HarmonicEval framework consists of three steps. (a) Criterion-wise scoring is performed by prompting a VLM to evaluate the input text based on each criterion. (b) Score smoothing improves robustness based on the first-order statistics of the output token probability distributions. (c) Score aggregation produces an overall score using harmonic weighting based on the second-order statistics.
Second, as discussed in previous studies Kasai et al. (2022); Fabbri et al. (2021), existing metrics can fail to consider specific evaluation criteria when predicting overall scores. For instance, criteria such as correctness and completeness might be prioritized while conciseness is overlooked, as illustrated in Figure 1 (b). This limitation can become more pronounced when handling a range of vision-language tasks beyond image captioning, where comprehensive evaluation is crucial to achieve better alignment with human judgments.
To address these limitations, we introduce HarmonicEval, a reference-free evaluation metric that aggregates criterion-wise scores to produce overall scores in a bottom-up manner. HarmonicEval can evaluate the quality of text across diverse vision-language tasks while providing numerical evaluation scores for each criterion, such as correctness, fluency, and conciseness. The evaluation procedure of HarmonicEval consists of three steps: 1) criterion-wise scoring, performed by prompting a VLM to evaluate text based on each specific criterion; 2) score smoothing, which enhances robustness by calculating the expected score from the output token probabilities assigned by the VLM; and 3) score aggregation, where the overall score is computed using a harmonic weighting scheme that places greater trust in criteria with lower variance, thus ensuring balanced and reliable assessments. By following these steps, HarmonicEval produces overall scores along with criterion-wise scores, offering a numerically explainable and comprehensive evaluation.
Furthermore, to promote research in this field, we provide the multi-task multi-criteria human evaluation (MMHE) dataset, a new human evaluation dataset spanning four vision-language tasks and five evaluation criteria. Specifically, MMHE consists of 18,000 human expert judgments for referring expression generation (REG), visual question answering (VQA), visual document understanding (VDU), and image captioning (IC). This dataset serves as a comprehensive benchmark for evaluating the alignment between automatic evaluation metrics and human judgments, which is essential for enhancing the reliability of VLMs.
Our experiments on MMHE demonstrate that HarmonicEval exhibits higher correlations with human judgments than conventional metrics, while providing criterion-wise scores. We also show that existing metrics indeed ignore certain evaluation criteria in some cases, resulting in less comprehensive evaluations. Furthermore, we evaluate HarmonicEval on the Flickr8k-EX and Flickr8k-CF Hodosh et al. (2013) datasets and show that HarmonicEval performs comparable to or better than state-of-the-art metrics.
2 Method
This section introduces HarmonicEval, a reference-free evaluation metric that integrates criterion-wise evaluation scores through harmonic weighting. Figure 2 shows an overview of HarmonicEval. First, a VLM is prompted to generate raw evaluation scores for the input text on each criterion independently (§2.1). Second, score smoothing is applied to obtain robust scores by leveraging the output token probabilities assigned by the VLM (§2.2). Finally, the individual scores are aggregated to produce an overall score through harmonic weighting (§2.3).
2.1 Prompts for criterion-wise evaluation
[…]
Correctness (Crt): The degree to which the target text accurately reflects the content of the input image and text.
Completeness (Cmp): The extent to which the target text captures all relevant and significant details of the input image and text.
Clarity (Clr): The ease with which the reader can understand the target text.
Fluency (Flu): The grammatical accuracy and natural flow of the target text.
Conciseness (Cnc): The efficiency of the target text in conveying information without unnecessary verbosity.
Table 1: Evaluation criteria.
The prompt $\bm{p}_{c}$ instructs the VLM to evaluate text with respect to the criterion $c$ on a five-point scale. Below is an example of the correctness criterion for the IC task.
Your task is to rate the candidate caption for the given image on a scale of 1 to 5 on the following criterion and rating scale.
Evaluation Criterion:
- Correctness: How accurately does the caption describe the image?
Rating Scale:
- 1 Very Low Correctness: The caption is mostly or entirely incorrect …
- 5 Extremely High Correctness: The caption perfectly captures all …
Here, the underlined portions indicate task-dependent variations in the content. For example, “caption” can be “answer” for the VQA task [^1] Details of evaluation and prompts are in Appendix C.1..
2.2 Score smoothing
To improve alignment with human judgment, this step applies score smoothing. Inspired by previous studies Liu et al. (2023b); Lee et al. (2024b), the first-order statistics of the output token probability distribution is utilized. Specifically, the smoothed scores $\tilde{s}_{c}$ are expected scores calculated as follows:
$\tilde{s}_{c}=\sum_{r=1}^{5}rp_{c}(r),$ (1)
where $p_{c}(r)$ is a probability of the token $r$ with respect to the criterion $c$ assigned by the VLM.
2.3 Score aggregation
Given a set of criterion-wise scores $\{\tilde{s}_{c}:c\in\mathcal{C}\}$, this step aggregates $\tilde{s}_{c}$ using harmonic weighting, which calculates the overall score $\bar{s}$ as
$\bar{s}=\sum_{c\in\mathcal{C}}w_{c}\tilde{s}_{c},$ (2)
where $0\leq w_{c}\leq 1$ are weighting coefficients. The weighting coefficients are determined based on the second-order statistics of $p_{c}$ as
$w_{c}=\frac{1}{H}\sigma_{c}^{-2(1-\gamma)/\gamma},$ (3)
where $\sigma_{c}$ is the standard deviation of the output token probabilities given by
$\sigma_{c}=\sqrt{\sum_{r=1}^{5}(r-\tilde{s}_{c})^{2}\ p_{c}(r)},$ (4)
and $H=\sum_{c}\sigma_{c}^{-2(1-\gamma)/\gamma}$ is the harmonic mean of the variances with a hyperparameter $\gamma$. Note that smaller values of $\sigma_{c}$ can be interpreted as indicating higher confidence in the evaluation of $c$.
The role of hyperparameter $\gamma$ is to bridge three weighting strategies: uniform weighting, inverse variance weighting and selective weighting. Below, we detail each strategy and discuss the hyperparameter setting.
Figure 3: MMHE dataset is a multi-task multi-criteria human evaluation dataset. Each candidate text is manually evaluated by three expert annotators.
Uniform weighting.
When $\gamma$$=$$1$, harmonic weighting reduces to uniform weighting $w_{c}$$=$$1/|\mathcal{C}|$. This is effective when all criterion-wise scores are equally reliable in determining overall scores. However, this may not be the best estimator as observed variances are ignored in the aggregation process.
Inverse variance weighting.
When $\gamma$$=$$0.5$, harmonic weighting reduces to inverse variance weighting $w_{c}\propto\sigma_{c}^{-2}$. This provides the best linear unbiased estimator under the assumption that the observed variance is due to statistical fluctuations. However, this assumption is not always reasonable, as each criterion may have its own variance.
Selective weighting.
When $\gamma\to 0$, only the score $\tilde{s}_{c}$ with the smallest variance is selected as the overall score. This approach is used in experiments to show the necessity of aggregation.
Discussion. Considering that the evaluation criteria are carefully designed, the uniform weighting ($\gamma=1.0$) may align closely with human expert judgment, and $1.00.5\leq\gamma<$ 1.0 may further improve the alignment because it reflects the confidence of criterion-wise scores. Since the assumption underlying the inverse variance weighting ($\gamma=0.5$) may not be reasonable, we hypothesize that a value between 0.5 and 1.0 is optimal and choose $\gamma=0.75$ as the default value. The detailed mathematical background is provided in Appendix A.
3 Dataset
This section presents the MMHE dataset, which consists of 18,000 human expert judgments covering four vision-language tasks and five evaluation criteria. This dataset promotes research towards automatic evaluation that aligns with human evaluation in a fine-grained manner. Example human judgment scores are shown in Figure 3.
3.1 Vision-language tasks
To assess the multifaceted capabilities of VLMs, we selected four vision-language tasks: REG, VQA, VDU, and IC.
Referring expression generation (REG)
aims to generate textual expressions that uniquely identify a specific object marked by a bounding box. To evaluate the expression quality, the VLM is given an expression and an image with the relevant bounding box highlighted in red.
Visual question answering (VQA)
aims to answer a question posed in natural language based on the content of an image. To evaluate the answer quality, the VLM is provided with a triplet consisting of an image, a question, and an answer.
Visual document understanding (VDU)
focuses on interpreting and extracting meaningful information from visually presented documents. Quality evaluation follows a question-answering manner. Specifically, the VLM is provided with a document image, along with a question-answer pair related to the content.
Image captioning (IC)
aims to generate captions for given images. To evaluate caption quality, the VLM is provided with a text-image pair.
3.2 Dataset construction
The dataset construction process consists of three steps: 1) Source selection, selecting input texts and images for tasks from datasets; 2) Target generation, creating target texts from sources using state-of-the-art VLMs; and 3) Human expert evaluation, assessing the quality of the generated texts.
Source selection.
We selected the following four datasets: RefCOCO Kazemzadeh et al. (2014) for REG, OK-VQA Marino et al. (2019) for VQA, VisualMRC Tanaka et al. (2021) for VDU, and MSCOCO Lin et al. (2014) for IC. We randomly sampled 100 instances from the validation or test subsets of each dataset.
Target generation.
The target texts to be evaluated were generated using state-of-the-art VLMs. Specifically, we employed ten VLMs: LLaVA-1.5-7B/13B Liu et al. (2023a), InstructBLIP-Vicuna-7B/13B Dai et al. (2023), Qwen-VL Bai et al. (2023), Qwen2-VL-Instruct-7B/72B Wang et al. (2024a), CogVLM-Chat Wang et al. (2024b), GPT-4o-mini and GPT-4o OpenAI (2024). These VLMs include both proprietary and open-source models, as well as various model sizes, which lead to variability in the generated target text. The details of prompts and hyperparameters used for each VLM are provided in Appendix C.2.
Human expert evaluation.
Five expert annotators were provided with an explanation of the vision-language tasks and instructed to carefully review the definitions of the evaluation criteria. They then rated each target text for each criterion on a five-point scale. After completing the evaluations, they conducted a thorough review to ensure consistency. Each instance was scored independently by three annotators, resulting in a total of 18,000 human judgment scores.
Overall judgments.
To collect overall judgments, we adopted a best-of-three approach, in which annotators are asked to choose the best text from three candidates. We did not use a five-point scale for the overall score because defining each rating level (from 1 to 5) for overall quality is difficult and could introduce bias Chiang et al. (2024).
Figure 4: Score distributions of the MMHE dataset.
3.3 Data Analysis
Figure 4 presents the score distributions for each criterion across the four tasks. The correctness and completeness criteria exhibit diverse score distributions for most tasks, suggesting that even the state-of-the-art VLMs face challenges in these aspects. In contrast, the fluency criterion shows a narrow score distribution across most tasks, with a dominant score of 5. This indicates that the VLMs generate fluent text even when visual understanding is inaccurate. Nonetheless, we consider fluency an essential criterion, as overlooking it could lead to high scores being assigned to correct but non-fluent text. A qualitative example illustrating this is provided in Section 4.3. The clarity and conciseness criteria have score distributions that are intermediate between the diverse and narrow distributions. While the score of 5 is more prevalent than in the diverse distributions, other scores still occur with noticeable frequency.
Method REG VQA VDU IC Avg.
Reference-based BLEU 45.3 29.4 57.3 46.8 44.7
ROUGE 49.0 30.8 56.0 47.9 45.9
CIDEr 42.5 25.0 62.1 42.7 43.1
METEOR 44.4 29.4 59.7 53.6 46.8
BERT-S 46.2 33.8 62.1 53.1 48.8
BART-S 56.4 20.5 60.9 57.8 48.9
Ref.-free CLIP-S 60.1 39.7 60.9 52.0 53.2
FLEUR 62.9 76.4 60.9 73.9 68.5
HarmonicEval (ours) 66.6 76.4 73.4 77.0 73.4
Table 2: Overall accuracy (%) on MMHE. The best result for each task is marked in bold. Average (Avg.) indicates the average performance across the four tasks.
Metric REG VQA VDU IC
Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc
BLEU $6.0$ $6.9$ $3.9$ $1.2$ $6.1$ $1.3$ $10.4$ $11.0$ $19.3$ $4.1$ $19.8$ $12.9$ $14.9$ $14.3$ $21.2$ $4.4$ $4.5$ $5.9$ $0.3$ $11.3$
ROUGE $2.3$ $5.7$ $4.4$ $3.5$ $3.9$ $7.1$ $2.8$ $5.0$ $8.1$ $10.2$ $20.0$ $14.7$ $16.2$ $17.9$ $22.7$ $5.2$ $6.5$ $9.0$ $4.4$ $9.7$
CIDEr $6.4$ $3.4$ $2.4$ $9.7$ $20.9$ $27.8$ $39.0$ $19.5$ $26.0$ $3.8$ $23.7$ $15.8$ $19.3$ $18.0$ $23.8$ $0.7$ $1.6$ $8.7$ $3.8$ $14.5$
METEOR $1.9$ $5.3$ $5.2$ $5.1$ $6.3$ $5.3$ $3.9$ $8.2$ $8.5$ $2.7$ $17.8$ $18.0$ $16.9$ $20.5$ $14.9$ $6.8$ $12.1$ $7.3$ $2.3$ $1.0$
BERT-S $6.5$ $6.9$ $6.5$ $8.6$ $12.4$ $2.8$ $14.3$ $4.9$ $10.0$ $6.1$ $21.0$ $17.4$ $20.4$ $21.6$ $23.9$ $12.3$ $11.1$ $6.4$ $4.7$ $10.5$
BART-S $4.4$ $6.7$ $4.2$ $7.8$ $3.1$ $13.4$ $20.2$ $2.8$ $16.6$ $1.6$ $22.4$ $21.3$ $21.6$ $17.9$ $14.7$ $4.8$ $4.3$ $4.3$ $2.2$ $3.2$
CLIP-S $13.5$ $14.4$ $6.8$ $0.9$ $5.1$ $6.6$ $5.4$ $7.2$ $8.1$ $4.5$ $15.2$ $12.5$ $15.0$ $12.6$ $8.4$ $20.2$ $21.3$ $11.1$ $3.2$ $3.5$
FLEUR
$\bm{29.3}$
$\bm{30.8}$ $18.6$ $8.7$ $11.2$ $38.7$ $38.2$
$\bm{39.9}$
$39.8$ $\bm{44.7}$ $38.1$ $37.1$ $44.6$ $35.2$ $28.2$ $33.9$ $35.0$
$\bm{25.9}$
$24.5$ $14.0$
HarmonicEval
23.2
$\bm{30.8}$
$\bm{24.0}$
$\bm{20.7}$
$\bm{23.8}$
$\bm{53.5}$
$\bm{50.6}$
31.8
$\bm{51.9}$
44.4
$\bm{60.0}$
$\bm{48.8}$
$\bm{47.9}$
$\bm{51.2}$
$\bm{45.8}$
$\bm{44.7}$
$\bm{50.3}$
19.8
$\bm{36.4}$
$\bm{22.8}$
Table 3: Criterion-wise correlations on MMHE. Scores for the most positively and negatively correlated criteria are marked with black and gray underlines, respectively. Crt: Correctness, Cmp: Completeness, Clr: Clarity, Flu: Fluency, Cnc: Conciseness.
4 Experiments
4.1 Settings
Baselines.
We implemented eight baseline metrics for the experiments on MMHE. Four of these are traditional n-gram-based metrics: BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), and CIDEr Vedantam et al. (2015). The remaining three metrics leverage neural network-based approaches: BERTScore Zhang et al. (2020), BARTScore Yuan et al. (2021), and CLIPScore Hessel et al. (2021). Additionally, we employ FLEUR Lee et al. (2024b), a VLM-based evaluation metric that provides a single overall score and its textual explanation. Details of these baselines are provided in Appendix E.
Performance measures.
To assess the performance of evaluation metrics on MMHE, we use accuracy (%) for overall evaluations and Kendall’s tau correlation coefficient $\tau$ for criterion-wise evaluations.
Implementation details.
The GPT-4o model is used as a backbone of HarmonicEval. The temperature hyperparameter is set to zero so that the outputs are as deterministic as possible for reproducibility. More details are in Appendix C.
4.2 Main results
Overall performance.
Table 2 compares HarmonicEval with eight metrics in terms of overall performance on MMHE. HarmonicEval achieves the highest accuracy on REG (66.6), VQA (76.4), VDU (73.4), and IC (77.0). While FLEUR also achieves the highest score of 76.4 on VQA, it falls short on VDU. These results underscore the effectiveness of HarmonicEval across multiple tasks.
Criterion-wise performance.
Table 3 shows the correlations between predicted scores and human judgments for each evaluation criterion. HarmonicEval evaluates numerical explainability by reporting correlations between its predicted criterion-wise scores and human judgment scores, whereas other metrics report correlations between their predicted overall scores and criterion-wise human judgment scores. As shown, HarmonicEval achieves the highest correlation across most criteria, highlighting its numerical explainability. FLEUR emerges as the second-best metric, a trend also observed in overall comparisons.
Task-wise trends in baseline metrics.
To gain deeper insights into the evaluation behavior of baseline metrics, we examine the task-wise correlations. Specifically, we explore which evaluation criteria baseline metrics tend to prioritize or deprioritize for each task, potentially hindering comprehensive evaluation. In Table 3, black underlines denote the most correlated criterion, while gray underlines indicate the least correlated criterion for each metric within a task.
For REG, completeness shows the highest correlation across most metrics. This is understandable, as REG requires explicit expressions to identify a unique object by distinguishing it from surrounding objects. For VQA, most metrics are more strongly correlated with conciseness but less so with completeness. This indicates that existing metrics deprioritize completeness, potentially leading to inaccurate evaluations of insufficient answers. A similar trend is observed in VDU, where completeness is also deprioritized by existing metrics. For IC, fluency exhibits low correlations for most metrics, suggesting a tendency to assign high scores even to texts lacking natural flow. A qualitative example illustrating this tendency is provided in Section 4.3.
4.3 Analysis.
Figure 5: Qualitative examples.
Qualitative examples.
We present a qualitative comparison between HarmonicEval and FLEUR. To generate textual explanations for HarmonicEval, we prompt the VLM with “Why? Tell me the reason.”, following Lee et al. (2024b).
Figure 5 illustrates two examples from the IC task. In Figure 5 (a), HarmonicEval detects severe fluency issues and reflects these in the overall score, whereas FLEUR fails to identify such deficiencies. This highlights that FLEUR may overlook specific evaluation criteria, such as fluency, as discussed in Section 1. In Figure 5 (b), HarmonicEval identifies errors in the caption more accurately than FLEUR, assigning a low correctness score and providing a clear, criterion-specific explanation. This suggests that HarmonicEval is able to offer not only numerical explainability but also textual explanation.
Metric REG VQA VDU IC Avg.
HarmonicEval $66.6$ $\bm{76.4}$ $73.4$ $\bm{77.0}$ $\bm{73.4}$
w/o criterion-wise scoring
$62.0$ $73.5$ $\bm{75.9}$ $76.5$ $72.0$
w/o harmonic weighting
$65.7$ $75.0$ $73.4$ $76.5$ $72.6$
w/o score smoothing
$\bm{67.5}$ $70.5$ $70.4$ $72.4$ $70.2$
Table 4: Ablation study. In the "w/o criterion-wise scoring" setting, we prompt the VLM to predict the overall score directly.
$\gamma$ REG VQA VDU IC Avg.
$0.01$ $47.2$ $69.1$ $61.4$ $53.1$ $57.7$
$0.50$ $\bm{66.6}$ $\bm{76.4}$ $\bm{73.4}$ $76.5$ $73.2$
$0.75$ $\bm{66.6}$ $\bm{76.4}$ $\bm{73.4}$ $\bm{77.0}$ $\bm{73.4}$
$1.00$ $65.7$ $75.0$ $\bm{73.4}$ $76.5$ $72.6$
Table 5:
Hyperparameter study.
VLM REG VQA VDU IC Avg.
LLaVA-7B
$62.9$ $72.0$ $61.4$ $73.4$ $67.4$
LLaVA-13B
$64.8$ $\bm{77.9}$ $63.2$ $72.9$ $69.7$
GPT-4o
$\bm{66.6}$ $76.4$ $\bm{73.4}$ $\bm{77.0}$ $\bm{73.4}$
Table 6: Overall accuracy (%) for different VLMs.
Ablation study.
Table 4 highlights the effectiveness of three key components in HarmonicEval: criterion-wise scoring, score smoothing, and harmonic weighting. As shown, omitting criterion-wise scoring reduces performance on REG, VQA, and IC, indicating that comprehensive evaluation improves overall accuracy. Similarly, omitting harmonic weighting or score smoothing lowers performance across most tasks, demonstrating that these techniques stabilize the evaluation process.
VLM REG VQA VDU IC
Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc
LLaVA-7B
$14.5$ $14.0$ $14.8$ $\bm{28.4}$ $2.9$ $42.2$ $49.4$ $23.7$ $45.2$ $18.1$ $29.7$ $40.3$ $28.6$ $44.9$ $7.2$ $28.8$ $31.5$ $7.8$ $34.8$ $10.5$
LLaVA-13B
$\bm{31.4}$ $28.2$ $22.4$ $26.1$ $15.0$ $49.4$ $\bm{57.8}$ $26.7$ $49.7$ $30.5$ $39.7$ $45.9$ $36.4$ $42.2$ $27.7$ $35.6$ $40.0$ $4.5$ $27.4$ $5.2$
GPT-4o $23.2$ $\bm{30.8}$ $\bm{24.0}$ $20.7$ $\bm{23.8}$ $\bm{53.5}$ $50.6$ $\bm{31.8}$ $\bm{51.9}$ $\bm{44.4}$ $\bm{60.0}$ $\bm{48.8}$ $\bm{47.9}$ $\bm{51.2}$ $\bm{45.8}$ $\bm{44.7}$ $\bm{50.3}$ $\bm{19.8}$ $\bm{36.4}$ $\bm{22.8}$
Table 7: Criterion-wise performance for different VLMs.
Hyperparameter study.
Table 5 shows the hyperparameter study for $\gamma$. As expected, harmonic weighting using $1.00.5\leq\gamma<$ 1.0 performs better than uniform weighting ($\gamma=1.0$). Mathematical discussion underlying assumptions of these weightings is provided in Appendix A.
Backbone VLMs.
Tables 6 and 7 compare the overall performance and criterion-wise performance of three backbone VLMs: LLaVA-1.5-7B, LLaVA-1.5-13B, and GPT-4o. As shown, LLaVA-1.5-13B consistently perform better than LLaVA-1.5-13B. GPT-4o outperforms both models but leaves room for improvement, underscoring the need for future research to enhance alignment with human evaluations.
Metric Flickr8k-EX Flickr8k-CF
Reference-based BLEU 30.8 16.9
ROUGE 32.3 19.9
METEOR 41.8 22.2
CIDEr 43.9 24.6
SPICE 44.9 24.4
BERT-S 39.2 22.8
BERT-S++ 46.7
TIGEr 49.3
ViLBERTS-F 50.1
FAIEr-4 52.6 35.4
RefCLIP-Score 53.0 36.4
Polos $\bm{56.4}$ 37.8
RefFLEUR 51.9 38.8
Reference-free UMIC 46.8
FAIEr-r 50.1 32.4
CLIP-S 51.5 34.4
InfoCLIP 32.6 23.5
InfoMetIC 54.2 36.3
InfoMetIC+ 55.5 36.6
FLEUR 53.0 38.6
HarmonicEval $53.1$ $\bm{39.2}$
Table 8: Comparison on existing human evaluation datasets for IC task.
Performance on other IC datasets.
To verify the generalizability of HarmonicEval’s effectiveness across different datasets, we additionally evaluated it on two human evaluation datasets: Flickr8k-EX and Flickr8k-CF. These datasets, widely used for image captioning, allow comparisons with a broader range of baselines, including learning-based metrics specifically designed for IC  [^2] We use Kendall’s tau-c and tau-b for Flickr8k-EX and Flickr8k-CF, respectively, following prior studies Hessel et al. (2021); Lee et al. (2024b) The details of additional baselines Anderson et al. (2016); Jiang et al. (2019); Lee et al. (2020); Hu et al. (2023); Wada et al. (2024) are in Appendix E..
As shown in Table 8, HarmonicEval achieves performance comparable to or even better than state-of-the-art metrics on the three datasets. While our main focus is on automatic multi-task, multi-criteria evaluation, these results highlight the robustness of HarmonicEval across multiple datasets.
5 Related Work
Automatic evaluation metrics.
Traditional metrics such as BLEU Papineni et al. (2002) and ROUGE Lin (2004) were originally proposed for automatic evaluation of machine translation and text summarization models. These metrics rely on n-gram overlap with reference texts. In recent years, large language models (LLMs) have been increasingly employed as evaluation metrics for various natural language generation (NLG) tasks Kocmi and Federmann (2023); Chiang and Lee (2023); Zheng et al. (2023); Song et al. (2024). For example, G-Eval Liu et al. (2023b) introduced a form-filling paradigm for criterion-based evaluation for text summarization and dialogue generation. Our HarmonicEval further extends this line of research with its capability of handling various vision-language tasks and provides a robust framework for aggregating criterion-wise scores based on the second-order statistics of token probability distributions assigned by VLMs.
For vision-language tasks, CIDEr Vedantam et al. (2015) measures the consensus between candidate and reference captions by weighting n-grams using TF-IDF. Recent metrics leverage VLMs to offer more flexible evaluation paradigms Lee et al. (2024a); Yu et al. (2024); Zhuang et al. (2024); Maeda et al. (2024); Chen et al. (2024). FLEUR Lee et al. (2024b), a state-of-the-art reference-free metric, provides textual explanations that underlie its overall scores. However, existing metrics primarily focus on providing overall evaluation, lacking numerical explainability and often overlooking specific criteria. HarmonicEval addresses these issues by offering criterion-wise scores alongside an overall score, enabling more interpretable and comprehensive evaluations.
Human evaluation datasets.
Several human evaluation datasets have been developed for NLG tasks. For instance, the WMT shared tasks Semenov et al. (2023) provide benchmarks for machine translation, TMU-GFM Yoshimura et al. (2020) for grammatical error correction, and SummEval Fabbri et al. (2021) for text summarization. In the vision-language domain, existing datasets such as Flickr8k Hodosh et al. (2013), PASCAL-50S Vedantam et al. (2015), Composite Aditya et al. (2018), THUMB Kasai et al. (2022), and Polaris Wada et al. (2024) target image captioning. In contrast, our MMHE dataset provides a multi-task, multi-criteria human evaluation resource, including sentences generated by several state-of-the-art VLMs across diverse vision-language tasks. This dataset enables more nuanced evaluations that capture a variety of criteria and tasks, advancing the field beyond the focus on captioning alone.
6 Conclusion
We introduced HarmonicEval and MMHE, a novel metric and dataset for automatic multi-criteria evaluation for various vision-language tasks, respectively. HarmonicEval is a reference-free evaluation metric that utilizes score variance to aggregate individual scores in a statistically reliable way. MMHE is a dataset covering four vision-language tasks and five criteria. Experimental results demonstrated the effectiveness of HarmonicEval on both MMHE and existing human evaluation datasets. We hope our study contributes to advancing research in automatic evaluation metrics.
Limitations
This section discusses limitations from five perspectives: theory, modality, criteria, dataset, and model.
Theory.
The proposed harmonic weighting provides a statistically reliable way to aggregate individual scores, as discussed in Appendix A, with justification for using $1.00.5<\gamma<$ 1.0 based on clarified assumptions underlying inverse variance weighting and uniform weighting. However, there is no guarantee that $\gamma=0.75$ is the optimal setting for any tasks and VLMs. Further, since score distributions are approximated by the output token probabilities, there is a possibility of unintended bias. Notably, evaluation bias in LLM-based evaluation metrics has been documented in several studies Zheng et al. (2023); Liu et al. (2023b); Ohi et al. (2024). As such, further research into evaluation bias in VLM-based evaluation metrics is essential for future work.
Modality.
This study focused primarily on evaluating text quality in vision-language tasks because most state-of-the-art VLMs output only text. This leaves image quality evaluation underexplored. Given that several recent image generation models, such as DALL-E 3, are integrated into conversational systems using VLMs, the automatic evaluation of both generated image and text quality would be a promising next step toward the development of more user-friendly systems.
Criteria.
We carefully selected five general criteria that are considered effective across various vision-language tasks as detailed in Appendix B. These criteria were useful for discussions spanning the four vision-language tasks in this study. To further expand research to include a greater number of criteria and tasks, analyzing the relationship between task- or domain-specific criteria and these general criteria would also be necessary in future work.
Dataset.
As the number of criteria increased, it became difficult even for experts to maintain annotation consistency, leading to greater time requirements for data collection. MMHE extracted one hundred images from each task, which was considered to be a statistically reliable number, and each target text was evaluated by three annotators. While large-scale crowdsourcing was attempted to scale up the dataset, obtaining human judgments that adhered accurately to each rating scale was challenging because careful explanation of the rating process through direct communication was required. This leaves scaling up the number of tasks and images challenging.
Model.
Improving VLMs to generate better text based on the evaluation results remains future work. In particular, achieving high scores across all criteria within learning frameworks such as in-context learning or reinforcement learning would be an intriguing direction for further exploration.
Ethics Statement
Data collection.
We created and publicly release a new dataset as part of this research. The data collection process was conducted with careful consideration for ethical guidelines. All annotators were informed about the purpose of this dataset and provided consent before participation. Any personally identifiable information has been removed to ensure privacy protection. The dataset was reviewed to minimize harmful content, offensive language, or biases that could negatively impact downstream applications. However, some inherent biases might still be present due to the nature of the data sources including natural images.
Reproducibility.
All code necessary to reproduce the experimental results is provided in the supplementary material and will be made publicly available. All experiments have been conducted as deterministically as possible by fixing random seeds and setting the temperature hyperparameters to zero.
Appendix
Appendix A Mathematical background
This section provides a detailed mathematical background of harmonic weighting.
A.1 Notations and setting
Let $\mathcal{C}$ be a finite discrete set and $\hat{z}_{c}~{}(c\in\mathcal{C})$ be independent estimators for the overall score $s$. Harmonic weighting assumes that each $c\in\mathcal{C}$ has an intrinsic variance $\varsigma_{c}^{2}$ and the empirical distributions of individual scores $\hat{s}_{c}=\varsigma_{c}(\hat{z}_{c}-s)+s,\hat{s}_{c}\sim p^{*}_{c}$ are observable, where $p^{*}_{c}:\mathbb{R}\to\mathbb{R}_{+}$ is a probability density function that can be approximated by the output token probability distributions. Then, the following combined estimator $\bar{s}$ is used to predict the overall score:
$\bar{s}=\sum_{c\in\mathcal{C}}w_{c}\hat{z}_{c},$ (5)
where $0\leq w_{c}\leq 1$ are weighting coefficients, satisfying $\sum_{c=1}^{C}w_{c}=1$.
A.2 Best linear unbiased estimator
Harmonic weighting chooses the weight coefficients such that the combined estimator is the best linear unbiased estimator when all estimators $\hat{s}_{c}$ are unbiased, and the variance $\mathbb{V}[\hat{s}_{c}]=\sigma_{c}^{2}$ is finite. Since the best linear unbiased estimator minimizes the variance $\mathbb{V}[\bar{s}]$, the weighting coefficients are obtained by solving the following minimization problem:
$\min\mathbb{V}[\bar{s}]\,\text{ subject to }\,\sum_{c\in\mathcal{C}}w_{c}=1.$ (6)
Here, since we have
$\mathbb{V}[\bar{s}]=\sum_{c\in\mathcal{C}}\frac{w_{c}^{2}\sigma_{c}^{2}}{\varsigma_{c}^{2}},$ (7)
the solution is given by
$w_{c}=\frac{\alpha_{c}}{\sum_{c\in\mathcal{C}}\alpha_{c}},\ \ \alpha_{c}=\frac{\varsigma_{c}^{2}}{\sigma_{c}^{2}}\ \ (\forall c\in\mathcal{C}).$ (8)
A.3 Inverse variance weighting
Inverse variance weighting is a statistical technique used to combine multiple independent estimates, which assigns larger weights to reliable estimates having smaller variances. We have Proposition 1 under Assumption 1.
Assumption 1. The observed variance is entirely due to statistical fluctuations, i.e., $\varsigma_{c}=1$.
Proposition 1. Harmonic weighting reduces to inverse variance weighting under Assumption 1.
The proof is trivial from Eq. (8). In practice, the distribution $p^{*}_{c}$ is approximated by the output token probability distribution $p_{c}$, thereby $\sigma_{c}$ is approximated as $\sigma_{c}\simeq\hat{\sigma}_{c}$ where $\hat{\sigma}_{c}$ is an observation given by
$\hat{\sigma}_{c}$ $=\sqrt{\sum_{r=1}^{5}(r-\tilde{s}_{c})^{2}\ p_{c}(r)},$ (9)
$\tilde{s}_{c}$ $=\sum_{r=1}^{5}rp_{c}(r).$ (10)
A.4 Uniform weighting
For uniform weighting, we have Proposition 2 under Assumption 2.
Assumption 2. The observed variance is entirely intrinsic to each $c\in\mathcal{C}$ i.e., $\varsigma_{c}\propto\sigma_{c}$.
Proposition 2. Harmonic weighting reduces to uniform weighting under Assumption 2.
Proof.
Since a constant $\kappa>0$ exists such that $\varsigma_{c}=\kappa\sigma_{c}$, we have
$\alpha_{c}=\frac{(\kappa\sigma_{c})^{2}}{\sigma_{c}^{2}}=\kappa^{2}.$ (11)
Therefore, we have uniform weighting coefficients:
$w_{c}=\frac{\kappa^{2}}{\sum_{c\in\mathcal{C}}\kappa^{2}}=\frac{1}{|\mathcal{C}|}.$ (12)
A.5 Harmonic weighting
Considering that both Assumptions 1 and 2 make strong assumptions, it is reasonable to consider an intermediate case where part of the observed variance is due to statistical fluctuations and part is intrinsic to each $c\in\mathcal{C}$. If this hypothesis is correct, it is optimal to estimate $\varsigma_{c}$ using a hyperparameter $\gamma$ such that it represents a balance between inverse variance weighting and uniform weighting. Harmonic weighting formulates this with linear weighting in a log scale as
$\log\varsigma_{c}^{2\gamma}=\gamma\log({\varsigma}^{(1)}_{c})^{2}+(1-\gamma)\log({\varsigma}^{(2)}_{c})^{2}$ (13)
where ${\varsigma}^{(1)}_{c}=1$ is from Assumption 1 and ${\varsigma}^{(2)}_{c}=\sigma_{c}$ is from Assumption 2. In summary, harmonic weighting adopts Assumption 3 and has Propositions 3 and 4.
Assumption 3. The observed variance is partially due to statistical fluctuations and the remaining portion is intrinsic to each $c\in\mathcal{C}$.
Proposition 3. Harmonic weighting approaches to inverse variance weighting when $\gamma\to 0.5$.
Proposition 4. Harmonic weighting approaches to uniform weighting when $\gamma\to 1$.
We believe Assumption 3 is reasonless, and this justifies why $1.00.5<\gamma<$ 1.0 is used in practice.
Appendix B Evaluation Criteria
We selected five key criteria essential for evaluating the output text of vision-language tasks: correctness, completeness, clarity, fluency, and conciseness.
Correctness and completeness assess the alignment between the task input (e.g., an image for image captioning (IC), or an image and question for visual question answering (VQA)) and the task output (e.g., a candidate caption for IC, or a candidate answer for VQA). As described in Table 1, correctness measures how accurately the text reflects the content of the input image and text, while completeness evaluates whether the text includes sufficient and relevant information. Prior works have adopted similar schemes, such as Composite Aditya et al. (2018), which uses correctness and thoroughness, and Thumb Kasai et al. (2022), which employs precision and recall for evaluating IC tasks.
The remaining three criteria, clarity, fluency, and conciseness, focus on the intrinsic quality of the text. Drawing on prior studies in natural language generation evaluation Asano et al. (2017); Kryscinski et al. (2019); Fabbri et al. (2021); Freitag et al. (2021); Song et al. (2024), we adopted these metrics. While coherence and grammar were considered, they were not included: coherence is more suitable for tasks involving long-form text, such as summarization, and grammar is more relevant for tasks emphasizing grammatical correctness, such as grammatical error correction.
To validate these five criteria, we manually examined 100 instances across various vision-language tasks and confirmed their sufficiency for evaluating output text.
Appendix C Implementation details
This section provides details of full prompts for HarmonicEval and vision-language models used in the present study.
C.1 Full prompts
Correctness (Crt).
The correctness is the degree to which the target text accurately reflects the content of the input image and text. The following is the full prompt designed for the correctness criterion.
Your task is to rate the candidate caption for the given image on a scale of 1 to 5 on the following criterion and rating scale.
Evaluation Criterion:
- Correctness: How accurately does the caption describe the image?
Rating Scale:
- 1 Very Low Correctness: The caption is mostly or entirely incorrect; it fails to accurately describe the image or misses key elements.
- 2 Low Correctness: The caption includes some correct information but misses major aspects or includes incorrect details.
- 3 Moderate Correctness: The caption is generally accurate, capturing a basic understanding of the image but may omit some details or include minor inaccuracies.
- 4 High Correctness: The caption is mostly accurate, including key elements and details of the image with minimal inaccuracies.
- 5 Extremely High Correctness: The caption perfectly captures all relevant and discernible details of the image with complete accuracy.
Completeness (Cmp).
The completeness is the extent to which the target text captures all relevant and significant details of the input image and text. The following is the full prompt designed for the completeness criterion.
Your task is to rate the candidate caption for the given image on a scale of 1 to 5. Output the evaluation score first based on the following criteria and rating scale.
Evaluation Criteria:
- Completeness: How completely does the caption cover all the relevant and essential aspects of the image?
Rating Scale:
- 1 Very Low Completeness: The caption includes almost none of the essential aspects of the image, missing most key elements and details.
- 2 Low Completeness: The caption mentions a few essential aspects but omits many key elements and relevant details.
- 3 Moderate Completeness: The caption covers some of the essential aspects of the image but misses several important details.
- 4 High Completeness: The caption includes most of the essential aspects of the image, with only a few minor elements missing.
- 5 Extremely High Completeness: The caption fully covers all relevant and essential aspects of the image, leaving no important details unaddressed.
Clarity (Clr).
The clarity is the ease with which the reader can understand the target text. The following is the full prompt designed for the comprehensiveness criterion.
Your task is to rate the text on a scale of 1 to 5. Output the evaluation score first based on the following criteria and rating scale.
Evaluation Criteria:
- Clarity: How clear is the text?
Rating Scale:
- 1 Ambiguous: The text is very unclear and can be interpreted in multiple ways, leading to significant confusion about its intended meaning.
- 2 Somewhat ambiguous: The text includes phrases or structures that make it non-definitive, such as using ’or’ or other ambiguous language, allowing for multiple interpretations.
- 3 Neutral: The text is moderately clear but may still leave some room for interpretation.
- 4 Clear: The text clearly describes the subject matter and is easily understood, with only minor imperfections.
- 5 Very Clear: The text is exceptionally clear, leaving no room for misinterpretation.
Fluency (Flu).
The fluency is the grammatical accuracy and natural flow of the target text. The following is the full prompt designed for the fluency criterion.
Your task is to rate the text on a scale of 1 to 5. Output the evaluation score first based on the following criteria and rating scale.
Evaluation Criteria:
- Fluency: How well the text is written in terms of grammar, punctuation, and phrasing.
Rating Scale:
- 1 Disfluent: The text contains numerous errors, making it difficult to understand.
- 2 Somewhat disfluent: The text has several noticeable errors that make it sound unnatural.
- 3 Moderately fluent: The text is generally understandable but contains errors that cause some discomfort while reading.
- 4 Fluent: The text flows well and is easy to understand with only minor imperfections.
- 5 Very fluent: The text is perfectly constructed with no grammatical errors or awkward phrasing.
Conciseness (Cnc).
The conciseness is the efficiency of the target text in conveying information without unnecessary verbosity. The following is the full prompt designed for the conciseness criterion.
Your task is to rate the text on a scale of 1 to 5. Output the evaluation score first based on the following criteria and rating scale.
Evaluation Criteria:
- Conciseness: How concise is the text?
Rating Scale:
- 1 Excessively verbose: The text is unnecessarily long-winded, containing redundant information or irrelevant details that significantly impair readability.
- 2 Overly verbose: The text includes unnecessary information or repetitive phrases that could be condensed without losing meaning.
- 3 Moderately verbose: The text is somewhat wordy but generally acceptable, with a few instances where information could be expressed more concisely.
- 4 Concise: The text is to-the-point and efficient, but there may be one or two phrases that could be slightly more condensed without losing meaning.
- 5 Optimally concise: The text conveys all necessary information using the minimum number of words possible, with no room for further condensation without losing meaning.
Here, the underlined portions indicate task-dependent variations in the content. For example, “caption"" can be “answer"" for the VQA task. In Table 8, only the correctness and comprehensiveness prompts are used because datasets contain only short captions and for computational efficiency. In Tables 6 and 7, the prompts are slightly modified for each VLM to ensure that only the score is returned as the response.
C.2 Prompts for target text generation
The following task-specific prompts are used to generate target texts.
Referring expression generation (REG).
To generate expressions for the target object within the corresponding bounding box, which is drawn in red in the input image, the following prompt is used:
Please carefully observe the area within the red bounding box in the image and come up with a caption for that area.
Visual question answering (VQA).
To generate answers for the VQA task, the following prompt is used:
[QUESTION] Please give an answer and explain it in two or three sentences.
where [QUESTION] is a placeholder for input questions.
Visual document understanding (VDU).
Because the VDU task is performed in a question-answering format, the same prompt as for VQA is used:
[QUESTION] Please give an answer and explain it in two or three sentences.
where [QUESTION] is a placeholder for input questions.
Image captioning (IC).
For caption generation, the following prompt is used:
Please provide a one-sentence caption for the provided image.
C.3 Vision language models
The target texts are generated by the following VLMs.
  • LLaVA-1.5-7B Liu et al. (2023a) is an open-source VLM based on the transformer architecture obtained by fine-tuning Vicuna on instruction data generated by GPT. The version is 1.5 and the number of parameters is 7 billion.
  • LLaVA-1.5-13B Liu et al. (2023a) is an extended version of LLaVA-1.5, featuring 13 billion parameters.
  • InstructBLIP-Vicuna-7B Dai et al. (2023) is an VLM that combines the BLIP-2 model with the Vicuna-7B language model by an instruction-aware Q-former module. This model is trained on a collection covering 11 tasks and 26 publicly available datasets.
  • InstructBLIP-Vicuna-13B Dai et al. (2023) is an extended version of InstructBLIP-Vicuna using Vicuna-13B.
  • Qwen-VL Bai et al. (2023) is an VLM that incorporates the vision transformer into the Qwen language model through cross-attention. This model is trained on a large-scale web-crawled dataset, including English and Chinese language subsets of LAION-5B.
  • Qwen2-VL-Instruct-7B/72B Wang et al. (2024a) is the instruction-tuned version of the Qwen2-VL model.
  • CogVLM-Chat Wang et al. (2024b) is a conversational VLM fine-tuned for dialogue-based tasks. This model is based on CogVLM-17B trained on publicly available datasets including LAION-2B and COYO-700M. Fine-tuning is performed on VQA and instruction-tuning datasets.
  • GPT-4o OpenAI (2024) is a VLM provided by OpenAI. Specifically, we used the model version gpt-4o-2024-08-06.
  • GPT-4o-mini is a lightweight variant of GPT-4o, optimized for efficiency and speed.
C.4 HarmonicEval
We use the GPT-4o model as the backbone of HarmonicEval. Specifically, we utilize gpt-4o-2024-08-06 as the model version in the API calls. For the criteria of Clarity, Fluency, and Conciseness, we input only the target text into the model, as these criteria are designed to assess the quality of the target text independently, without considering the source or the image. All experiments are conducted in a single attempt.
Appendix D MMHE
This section provides details of MMHE.
D.1 Source data
We select input texts and images for tasks from existing datasets. English is the language used for all datasets.
Referring expression generation (REG).
We adopt RefCOCO (Apache-2.0 license) Kazemzadeh et al. (2014) as a source dataset. This dataset contains 142,209 referring expressions for 50,000 objects in 19,994 images. Our usage of the data complies with the terms of its license.
Visual question answering (VQA).
We adopt OK-VQA (Apache-2.0 license) Marino et al. (2019) as a source dataset. This dataset contains 14,055 open-ended questions. Our usage of the data complies with the terms of its license.
Visual document understanding (VDU).
We adopt VisualMRC Tanaka et al. (2021) as a source dataset. This dataset contains over 30,000 question-answer pairs associated with more than 10,000 document images. Our usage of the data complies with the terms defined by its license.
Image Captioning (IC).
We adopt MSCOCO (CC BY 4.0) Lin et al. (2014) as a source dataset. Specifically, we utilized the 2017 Val split, which contains approximately 5,000 images and captions. Our usage of the data complies with the terms of its license.
D.2 Human Annotation
Five annotators who are English native speakers and experts in proofreading were hired for this process. The instructions provided to the annotators were almost identical to the prompts given to the VLM in HarmonicEval. On average, the annotators were paid \$20 per hour. Annotators were instructed on how the annotated data would be utilized.
D.3 Overall judgment
When calculating the overall judgment accuracy in Table 2, 4, and 6, we extract and utilize pairs of judgments from best-of-three triplets.
Appendix E Baselines
Below are details of the eight baselines implemented on the MMHE dataset.
  • Bilingual evaluation understudy (BLEU)  Papineni et al. (2002) is a metric for machine translation evaluation that measures the n-gram overlap between candidate and reference texts.
  • Recall-oriented understudy for gisting evaluation (ROUGE) Lin (2004) is a metric for text summarization and machine translation evaluation that measures the overlap of n-grams, skip-bigrams, and weighted longest common subsequences between candidate and reference texts.
  • Metric for evaluation of translation with explicit ordering (METEOR) Banerjee and Lavie (2005) is a metric for machine translation evaluation that assesses translations by explicit word-to-word matches between candidates and references.
  • Consensus-based image description evaluation (CIDEr) Vedantam et al. (2015) is a metric for image captioning that measures the consensus between candidate and reference captions by weighting n-grams based on their term frequency-inverse document frequency (TF-IDF).
  • BERTScore (BERT-S) Zhang et al. (2020) is a metric for text generation evaluation that measures the similarity between candidate and reference sentences using contextual embeddings extracted from BERT.
  • BARTScore (BART-S) Yuan et al. (2021) is a metric for text generation evaluation that computes weighted log probability using the BART model.
  • CLIPScore Hessel et al. (2021) is a metric for image captioning evaluation that measures the alignment between images and texts using CLIP embeddings.
  • FLEUR Lee et al. (2024b) is a state-of-the-art VLM-based metric that uses a two-level score smoothing technique. This metric utilizes LLaVa-1.5-13B. We also attempted to extend FLEUR for GPT-4o, but because the tokenizer of GPT-4o often outputs multi-digit numbers as a single token, it could not be implemented straightforwardly.
Below are details of the additional baselines for the Flickr8k-EX/CF and Composite datasets.
  • SPICE Anderson et al. (2016) is a reference-based metric for image captioning that compares semantic content using scene graphs.
  • TIGEr Jiang et al. (2019) is a reference-based metric for image captioning using the stacked cross attention neural network. It considers both how accurately a caption represents the image content and how closely machine-generated captions align with human-generated captions.
  • ViLBERTS-F Lee et al. (2020) is a reference-based metric for image captioning that computes pairwise cosine similarity between reference and candidate captions using the ViLBERT model.
  • FAIEr-4/r Wang et al. (2021) is a reference- and learning-based metric for image captioning evaluation. It introduces an attention module to fuse scene graph embeddings and produces scores based on fidelity and adequacy.
  • RefCLIP-S Hessel et al. (2021) is a reference-based variant of CLIPScore. It computes a harmonic mean of CLIPScore and the maximal reference cosine similarity.
  • Polos Wada et al. (2024) is a reference- and learning-based metric for image captioning evaluation. It combines CILP and RoBERTa via multimodal metric learning.
  • RefFLEUR Lee et al. (2024b) is a reference-based variant of FLEUR. It incorporates reference captions and a candidate caption into its prompt.
  • UMIC Lee et al. (2021) is a reference-free metric, trained to distinguish negative captions using contrastive learning.
  • InfoCLIP Hu et al. (2023) is a simplified version of InfoMetIC, which removes the intra and inter modality fusion as well as parameters in fine-grained scoring.
  • InfoMetIC Hu et al. (2023) is a reference-free learning-based metric for image captioning evaluation. It automatically constructs training examples from captions to produce detailed token-level scores.
Appendix F Computational Budget
We run all the experiments on ABCI (https://abci.ai/), Compute Node(A), whose CPUs are two Intel Xeon Platinum 8360Y, and GPUs are eight NVIDIA A100 SXM4. The approximate total processing time is 40 hours.

### v2

@@
-HarmonicEval: Multi-modal, Multi-task, Multi-criteria Automatic Evaluation Using a Vision Language Model
+Multi-modal, Multi-task, Multi-criteria Automatic Evaluation with Vision Language Models
 Abstract
-Vision-language models (VLMs) have shown impressive abilities in text and image understanding.
-However, existing metrics for evaluating the text generated by VLMs focus exclusively on overall quality, leading to two limitations: 1) it is challenging to identify which aspects of the text need improvement from the overall score; 2) metrics may overlook specific evaluation criteria when predicting an overall score.
-To address these limitations, we propose HarmonicEval, a reference-free evaluation metric that aggregates criterion-wise scores to produce the overall score in a bottom-up manner.
-Furthermore, we construct the Multi-task Multi-criteria Human Evaluation (MMHE) dataset, which comprises 18,000 expert human judgments across four vision-language tasks.
-Our experiments demonstrate that HarmonicEval achieves higher correlations with human judgments than conventional metrics while providing numerical scores for each criterion.
+Vision-language models (VLMs) have shown impressive abilities across a range of multi-modal tasks.
+However, existing metrics for evaluating the quality of text generated by VLMs typically focus on an overall evaluation for a specific task, such as image captioning.
+While the overall evaluation is essential for any task, the criteria prioritized can differ depending on the task, making it challenging for current metrics to adapt to multi-task scenarios.
+To address this limitation, we propose HarmonicEval, a reference-free comprehensive evaluation metric that aggregates criterion-wise scores to produce the overall score in a bottom-up manner.
+Furthermore, we construct the Multi-task Multi-criteria Human Evaluation (MMHE) dataset, which comprises 18,000 expert human judgments across four multi-modal tasks.
+Our experiments demonstrate that HarmonicEval achieves higher correlations with human judgments than conventional metrics while providing numerical scores for each criterion [^1] Our code and data will be available publicly..
 1 Introduction
-Automatic evaluation of text generated by vision-language models (VLMs) is essential for enhancing their robustness and reliability in tasks such as image captioning and visual question answering Hessel et al.
+Automatic evaluation of text generated by vision-language models (VLMs) is essential for improving their performance across various multi-modal tasks, such as image captioning and visual question answering Hessel et al.
 (2021); Lee et al.
 (2024b).
-As VLMs advance rapidly, the need for more informative and accurate metrics aligned with human perception has grown.
-However, most existing metrics focus solely on measuring the overall quality of texts generated by VLMs, lacking two critical aspects expected in automatic evaluation: 1) numerical explainability and 2) comprehensiveness.
-Figure 1: Necessity of multi-criteria evaluation.
-(a) Example of text quality evaluation in referring expression generation.
-HarmonicEval provides scores for each criterion.
-(b) Example in visual question answering where existing interpretable approach overlooks conciseness.
-First, evaluation metrics that provide only overall scores Papineni et al.
+As the range of tasks that VLMs can perform continues to expand, developing specialized evaluation metrics for each task becomes increasingly difficult.
+Hence, a comprehensive metric capable of evaluating text across multiple tasks is highly desirable.
+However, most existing metrics focus on measuring the overall quality of text within a specific task, limiting their applicability in multi-task settings (see Figure 1 (a)).
+Figure 1: Multi-task and multi-criteria evaluation.
+(a) Conventional single-criterion approach focuses on a single task, such as image captioning.
+(b) HarmonicEval integrates multiple criteria to provide overall scores.
+MMHE dataset consists of 18,000 expert human judgments across four multi-modal tasks and five criteria.
+Existing metrics that provide only overall scores Papineni et al.
 (2002); Zhang et al.
 (2020); Hessel et al.
-(2021) cannot identify specific deficiencies in aspects of the text, such as clarity or fluency, as illustrated in Figure 1(a).
-While recent metrics for vision-language tasks Chan et al.
-(2023); Lee et al.
-(2024b) offer free-form textual explanations to complement scores, these explanations lack the granularity and consistency that come from explicit numerical breakdowns.
-As a result, providing a numerical explanation that links specific aspect qualities to overall scores remains a challenge in the automatic evaluation of vision-language tasks.
-Figure 2: HarmonicEval framework consists of three steps.
-(a) Criterion-wise scoring is performed by prompting a VLM to evaluate the input text based on each criterion.
-(b) Score smoothing improves robustness based on the first-order statistics of the output token probability distributions.
-(c) Score aggregation produces an overall score using harmonic weighting based on the second-order statistics.
-Second, as discussed in previous studies Kasai et al.
+(2021) often prioritize specific evaluation criteria, as discussed in previous studies Kasai et al.
 (2022); Fabbri et al.
-(2021), existing metrics can fail to consider specific evaluation criteria when predicting overall scores.
-For instance, criteria such as correctness and completeness might be prioritized while conciseness is overlooked, as illustrated in Figure 1 (b).
-This limitation can become more pronounced when handling a range of vision-language tasks beyond image captioning, where comprehensive evaluation is crucial to achieve better alignment with human judgments.
-To address these limitations, we introduce HarmonicEval, a reference-free evaluation metric that aggregates criterion-wise scores to produce overall scores in a bottom-up manner.
-HarmonicEval can evaluate the quality of text across diverse vision-language tasks while providing numerical evaluation scores for each criterion, such as correctness, fluency, and conciseness.
-The evaluation procedure of HarmonicEval consists of three steps: 1) criterion-wise scoring, performed by prompting a VLM to evaluate text based on each specific criterion; 2) score smoothing, which enhances robustness by calculating the expected score from the output token probabilities assigned by the VLM; and 3) score aggregation, where the overall score is computed using a harmonic weighting scheme that places greater trust in criteria with lower variance, thus ensuring balanced and reliable assessments.
-By following these steps, HarmonicEval produces overall scores along with criterion-wise scores, offering a numerically explainable and comprehensive evaluation.
-Furthermore, to promote research in this field, we provide the multi-task multi-criteria human evaluation (MMHE) dataset, a new human evaluation dataset spanning four vision-language tasks and five evaluation criteria.
-Specifically, MMHE consists of 18,000 human expert judgments for referring expression generation (REG), visual question answering (VQA), visual document understanding (VDU), and image captioning (IC).
-This dataset serves as a comprehensive benchmark for evaluating the alignment between automatic evaluation metrics and human judgments, which is essential for enhancing the reliability of VLMs.
-Our experiments on MMHE demonstrate that HarmonicEval exhibits higher correlations with human judgments than conventional metrics, while providing criterion-wise scores.
-We also show that existing metrics indeed ignore certain evaluation criteria in some cases, resulting in less comprehensive evaluations.
-Furthermore, we evaluate HarmonicEval on the Flickr8k-EX and Flickr8k-CF Hodosh et al.
-(2013) datasets and show that HarmonicEval performs comparable to or better than state-of-the-art metrics.
-2 Method
-This section introduces HarmonicEval, a reference-free evaluation metric that integrates criterion-wise evaluation scores through harmonic weighting.
-Figure 2 shows an overview of HarmonicEval.
-First, a VLM is prompted to generate raw evaluation scores for the input text on each criterion independently (§2.1).
-Second, score smoothing is applied to obtain robust scores by leveraging the output token probabilities assigned by the VLM (§2.2).
-Finally, the individual scores are aggregated to produce an overall score through harmonic weighting (§2.3).
-2.1 Prompts for criterion-wise evaluation
+(2021).
+For example, metrics for evaluating image captions typically prioritize correctness and completeness over conciseness and fluency.
+When these metrics are applied to other tasks, such as visual question answering, they tend to overvalue verbose or unnatural responses.
+To address this limitation, integrating multiple evaluation criteria to predict the overall score, a concept we refer to as comprehensive evaluation, holds significant potential for a more comprehensive assessment in multi-task scenarios.
+However, this approach remains underexplored due to the lack of meta-evaluation datasets that provide human judgment annotations across multiple tasks and criteria.
+This motivates us to introduce a novel evaluation metric with a new dataset in this paper.
+Figure 2: HarmonicEval framework consists of two steps.
+(a) Criterion-wise scoring is performed by prompting a VLM to evaluate the input text based on each criterion, followed by score smoothing to improve robustness based on the first-order statistics.
+(b) Score aggregation produces an overall score using harmonic weighting based on the second-order statistics, aiming to reduce statistical fluctuations.
+First, we propose HarmonicEval, a reference-free harmonic evaluation metric for multiple multi-modal tasks.
+As shown in Figure 1 (b), HarmonicEval integrates multiple criteria to produce the overall score in a bottom-up manner.
+Specifically, the evaluation pipeline consists of two steps: 1) criterion-wise scoring, where a VLM is prompted to evaluate the input text based on each specific criterion, and 2) score aggregation, where the overall score is calculated from the criterion-wise scores.
+For score aggregation, we introduce a novel harmonic weighting scheme that automatically determines weight coefficients based on the second-order statistics of the output token probability distributions.
+Second, we introduce the multi-task multi-criteria human evaluation (MMHE) dataset, the first meta-evaluation dataset that provides human judgment annotations across multiple criteria and multiple multi-modal tasks.
+Specifically, MMHE consists of 18,000 expert human judgments on five criteria for four diverse tasks: referring expression generation (REG), visual question answering (VQA), visual document understanding (VDU), and image captioning (IC).
+This dataset serves as a benchmark for assessing the generalizability of automatic evaluation metrics.
+Our experiments on MMHE show that HarmonicEval achieves higher correlations with human judgments than conventional metrics, while also providing criterion-specific scores that highlight areas for improvement.
+Furthermore, we demonstrate that HarmonicEval achieves state-of-the-art or comparable performance in conventional image caption evaluation scenarios across five widely used datasets that provide only overall judgments: Flickr8k-EX / CF Hodosh et al.
+(2013), Composite Aditya et al.
+(2018), PASCAL-50S Vedantam et al.
+(2015), and FOIL Shekhar et al.
+(2017).
+In summary, our key contributions are threefold:
+  • We propose HarmonicEval, a novel reference-free metric for harmonic evaluation across multiple multi-modal tasks.
+  • We introduce MMHE, the first multi-task multi-criteria human evaluation dataset, consisting of 18,000 expert judgments spanning four multi-modal tasks and five evaluation criteria.
+  • We demonstrate the effectiveness of HarmonicEval on MMHE and five conventional image captioning datasets.
+In addition, we conduct the first in-depth analysis of how existing metrics implicitly prioritize different evaluation criteria.
+2 HarmonicEval
+As shown in Figure 2, the pipeline of HarmonicEval consists of two steps: criterion-wise scoring (§2.1) and score aggregation (§2.2).
+2.1 Criterion-wise scoring
+In this step, a VLM is employed as an evaluator and prompted to generate evaluation scores on each criterion independently.
+Let $\bm{t}$ be an input text to be evaluated, such as an image caption for the IC task.
+The evaluation process to obtain criterion-wise scores $s_{c}$ is formulated as $s_{c}=f([\bm{p}_{c},\bm{t}],\bm{x})$, where $c$ is a criterion, $\bm{x}$ is an input image, $f$ is a VLM, $\bm{p}_{c}$ is a prompt, and [,][,][ , ] denotes textual concatenation.
+To improve alignment with human judgements, score smoothing Liu et al.
+(2023b); Lee et al.
+(2024b) is applied as $\tilde{s}_{c}=\sum_{r\in R}rP(r\mid[\bm{p}_{c},\bm{t}],\bm{x})$ where $P(r\mid[\bm{p}_{c},\bm{t}],\bm{x})$ is the output token probability of for token $r$ assigned by the VLM, and $R$ is a set of ratings.
+2.2 Score aggregation
+To aggregate the criterion-wise scores $\tilde{s}_{c}$, we introduce harmonic weighting, a novel approach that leverages the second-order statistics of the output token probability distributions to adaptively determine the weight coefficients for aggregation.
+Compared to simple averaging, our aggregation approach aims to better align with human evaluation by dynamically emphasizing more reliable scores based on the input.
+Specifically, the overall score $S$ is computed as
+$S=\sum_{c\in\mathcal{C}}w_{c}\tilde{s}_{c},\quad w_{c}=\frac{1}{H}\sigma_{c}^{-2(1-\gamma)/\gamma},$ (1)
+where $w_{c}$ is a weight coefficient, $\mathcal{C}$ is a set of criteria, $\sigma_{c}$ is the standard deviation of the criterion-wise score given by
+$\sigma_{c}=\sqrt{\sum_{r\in R}(r-\tilde{s}_{c})^{2}\ P(r\mid[\bm{p}_{c},\bm{t}],\bm{x})},$ (2)
+and $H=\sum_{c\in\mathcal{C}}\sigma_{c}^{-2(1-\gamma)/\gamma}$ is the harmonic mean of the variances with a hyperparameter $\gamma$.
+Smaller values of $\sigma_{c}$ can be interpreted as indicating higher confidence in the evaluation of $c$.
+The role of hyperparameter $\gamma$ is to bridge three weighting strategies: uniform weighting, inverse variance weighting and selective weighting as detailed below.
+Uniform weighting.
+When $\gamma$$=$$1$, harmonic weighting reduces to uniform weighting $w_{c}$$=$$1/|\mathcal{C}|$.
+This is effective when all criterion-wise scores are equally reliable in determining the overall score.
+However, this does not provide the best estimator as observed variances are ignored in aggregation.
+Inverse variance weighting.
+When $\gamma$$=$$0.5$, harmonic weighting reduces to inverse variance weighting $w_{c}\propto\sigma_{c}^{-2}$.
+This provides the best linear unbiased estimator under the assumption that the observed variance is due to statistical fluctuations.
+Selective weighting.
+When $\gamma\to 0$, only the score $\tilde{s}_{c}$ with the smallest variance is selected as the overall score. This approach is used in experiments to show the necessity of aggregation.
+Discussion.
+When the evaluation criteria are carefully designed, the uniform weighting ($\gamma=1.0$) aligns closely with human expert judgment, and $1.00.5\leq\gamma<$ 1.0 further improves the alignment because it adaptively reflects the confidence of criterion-wise scores.
+Since the assumption underlying the inverse variance weighting ($\gamma=0.5$) is not reasonable when each criterion has its own variance, we hypothesize that a value between 0.5 and 1.0 is optimal and choose $\gamma=0.75$ as the default value.
+The detailed mathematical background is provided in Appendix A.
+2.3 Implementation details
 […]
+Aggregating these five criteria contributes to the overall text quality evaluation across a wide range of tasks.
+Depending on the task and the style of the input text, some criteria may not be necessary.
+However, when such evaluations are conducted, the output becomes less confident, leading to higher variance $\sigma_{c}$ and lower weight coefficients $w_{c}$ as VLMs account for task and criterion features in addition to the input text.
+Thus, HarmonicEval can adaptively perform comprehensive evaluations without manual tuning of weight coefficients.
+More details on criteria selection are in Appendix B.
 Correctness (Crt): The degree to which the target text accurately reflects the content of the input image and text.
@@
 Conciseness (Cnc): The efficiency of the target text in conveying information without unnecessary verbosity.
-Table 1: Evaluation criteria.
-The prompt $\bm{p}_{c}$ instructs the VLM to evaluate text with respect to the criterion $c$ on a five-point scale. Below is an example of the correctness criterion for the IC task.
-Your task is to rate the candidate caption for the given image on a scale of 1 to 5 on the following criterion and rating scale.
+Table 1: Five criteria for HarmonicEval and MMHE.
+Figure 3: MMHE dataset is a multi-task multi-criteria human evaluation dataset. Each candidate text is manually evaluated by three expert annotators.
+Prompts.
+The prompt $\bm{p}_{c}$ instructs the VLM to evaluate text with respect to the criterion $c$ on a five-point scale $R=\{1,2,3,4,5\}$.
+Below is the prompt of the correctness criterion for the IC task.
+Your task is to rate the caption for the given image on a scale of 1 to 5 on the following criterion and rating scale.
 Evaluation Criterion:
@@
 - 5 Extremely High Correctness: The caption perfectly captures all …
-Here, the underlined portions indicate task-dependent variations in the content. For example, “caption” can be “answer” for the VQA task [^1] Details of evaluation and prompts are in Appendix C.1..
-2.2 Score smoothing
-To improve alignment with human judgment, this step applies score smoothing.
-Inspired by previous studies Liu et al.
-(2023b); Lee et al.
-(2024b), the first-order statistics of the output token probability distribution is utilized.
-Specifically, the smoothed scores $\tilde{s}_{c}$ are expected scores calculated as follows:
-$\tilde{s}_{c}=\sum_{r=1}^{5}rp_{c}(r),$ (1)
-where $p_{c}(r)$ is a probability of the token $r$ with respect to the criterion $c$ assigned by the VLM.
-2.3 Score aggregation
-Given a set of criterion-wise scores $\{\tilde{s}_{c}:c\in\mathcal{C}\}$, this step aggregates $\tilde{s}_{c}$ using harmonic weighting, which calculates the overall score $\bar{s}$ as
-$\bar{s}=\sum_{c\in\mathcal{C}}w_{c}\tilde{s}_{c},$ (2)
-where $0\leq w_{c}\leq 1$ are weighting coefficients. The weighting coefficients are determined based on the second-order statistics of $p_{c}$ as
-$w_{c}=\frac{1}{H}\sigma_{c}^{-2(1-\gamma)/\gamma},$ (3)
-where $\sigma_{c}$ is the standard deviation of the output token probabilities given by
-$\sigma_{c}=\sqrt{\sum_{r=1}^{5}(r-\tilde{s}_{c})^{2}\ p_{c}(r)},$ (4)
-and $H=\sum_{c}\sigma_{c}^{-2(1-\gamma)/\gamma}$ is the harmonic mean of the variances with a hyperparameter $\gamma$.
-Note that smaller values of $\sigma_{c}$ can be interpreted as indicating higher confidence in the evaluation of $c$.
-The role of hyperparameter $\gamma$ is to bridge three weighting strategies: uniform weighting, inverse variance weighting and selective weighting.
-Below, we detail each strategy and discuss the hyperparameter setting.
-Figure 3: MMHE dataset is a multi-task multi-criteria human evaluation dataset. Each candidate text is manually evaluated by three expert annotators.
-Uniform weighting.
-When $\gamma$$=$$1$, harmonic weighting reduces to uniform weighting $w_{c}$$=$$1/|\mathcal{C}|$.
-This is effective when all criterion-wise scores are equally reliable in determining overall scores.
-However, this may not be the best estimator as observed variances are ignored in the aggregation process.
-Inverse variance weighting.
-When $\gamma$$=$$0.5$, harmonic weighting reduces to inverse variance weighting $w_{c}\propto\sigma_{c}^{-2}$.
-This provides the best linear unbiased estimator under the assumption that the observed variance is due to statistical fluctuations.
-However, this assumption is not always reasonable, as each criterion may have its own variance.
-Selective weighting.
-When $\gamma\to 0$, only the score $\tilde{s}_{c}$ with the smallest variance is selected as the overall score. This approach is used in experiments to show the necessity of aggregation.
-Discussion.
-Considering that the evaluation criteria are carefully designed, the uniform weighting ($\gamma=1.0$) may align closely with human expert judgment, and $1.00.5\leq\gamma<$ 1.0 may further improve the alignment because it reflects the confidence of criterion-wise scores.
-Since the assumption underlying the inverse variance weighting ($\gamma=0.5$) may not be reasonable, we hypothesize that a value between 0.5 and 1.0 is optimal and choose $\gamma=0.75$ as the default value.
-The detailed mathematical background is provided in Appendix A.
-3 Dataset
-This section presents the MMHE dataset, which consists of 18,000 human expert judgments covering four vision-language tasks and five evaluation criteria.
-This dataset promotes research towards automatic evaluation that aligns with human evaluation in a fine-grained manner.
+Here, the underlined portions indicate the task-dependent phrases. For example, “caption” is replaced with “answer” for the VQA task. The full prompts are provided in Appendix C.1.
+3 MMHE Dataset
+We present the MMHE dataset, the first meta-evaluation dataset that covers multiple evaluation criteria across multiple multi-modal tasks.
+We collected 18,000 expert human judgments spanning four tasks and five criteria.
 Example human judgment scores are shown in Figure 3.
-3.1 Vision-language tasks
-To assess the multifaceted capabilities of VLMs, we selected four vision-language tasks: REG, VQA, VDU, and IC.
-Referring expression generation (REG)
-aims to generate textual expressions that uniquely identify a specific object marked by a bounding box.
-To evaluate the expression quality, the VLM is given an expression and an image with the relevant bounding box highlighted in red.
-Visual question answering (VQA)
-aims to answer a question posed in natural language based on the content of an image.
-To evaluate the answer quality, the VLM is provided with a triplet consisting of an image, a question, and an answer.
-Visual document understanding (VDU)
-focuses on interpreting and extracting meaningful information from visually presented documents.
-Quality evaluation follows a question-answering manner.
-Specifically, the VLM is provided with a document image, along with a question-answer pair related to the content.
-Image captioning (IC)
-aims to generate captions for given images. To evaluate caption quality, the VLM is provided with a text-image pair.
+3.1 Dataset design
+Motivation.
+Despite numerous human evaluation datasets for multi-modal tasks, most focus on a single task (e.g., image captioning) or offer only an overall quality label.
+The purpose of MMHE is twofold: 1) to assess how evaluation metrics perform in a multi-task setting, and 2) to verify the hypothesis that each task may prioritize different criteria (e.g., correctness, completeness).
+To achieve these goals, we design MMHE to cover multiple multi-modal tasks and to explicitly collect human judgments across five evaluation criteria.
+Multi-modal tasks.
+We select four diverse tasks to illustrate how different criteria might matter in different contexts: 1) REG aims to generate a textual expression that uniquely identifies a specific object in the image (marked by a bounding box).
+We expect that completeness and clarity are crucial for precisely distinguishing the target object.
+2) VQA requires generating an answer to a question about the image content.
+Due to the nature of answering questions, we hypothesize that correctness and conciseness may be especially important here.
+3) VDU focuses on interpreting and extracting information from visually presented documents.
+Similar to VQA, we suspect that correctness and conciseness play key roles.
+4) IC involves producing a descriptive sentence for the entire image.
+We anticipate that correctness and completeness are especially relevant for capturing key elements.
+By spanning these tasks within a single dataset, MMHE enables a fine-grained, criterion-wise analysis of how metrics handle varied requirements.
+It is worth noting that MMHE is the first dataset to unify multiple multi-modal tasks and multiple evaluation criteria in a single benchmark.
 3.2 Dataset construction
-The dataset construction process consists of three steps: 1) Source selection, selecting input texts and images for tasks from datasets; 2) Target generation, creating target texts from sources using state-of-the-art VLMs; and 3) Human expert evaluation, assessing the quality of the generated texts.
+The dataset construction process consists of three steps: 1) Source selection, which selects source text-image pairs for each task; 2) Target generation, which creates target texts to be evaluated using state-of-the-art VLMs; and 3) Human expert evaluation, which assesses the quality of the target texts.
 Source selection.
@@
 (2014) for IC.
-We randomly sampled 100 instances from the validation or test subsets of each dataset.
+We randomly sampled 100 instances from the validation or test subset of each dataset.
 Target generation.
@@
 (2024b), GPT-4o-mini and GPT-4o OpenAI (2024).
-These VLMs include both proprietary and open-source models, as well as various model sizes, which lead to variability in the generated target text.
+These VLMs include both proprietary and open-source models, as well as various model sizes, which lead to variability in the generated target texts.
 The details of prompts and hyperparameters used for each VLM are provided in Appendix C.2.
 Human expert evaluation.
-Five expert annotators were provided with an explanation of the vision-language tasks and instructed to carefully review the definitions of the evaluation criteria.
+Five expert annotators were provided with an explanation of the four multi-modal tasks and instructed to carefully review the definitions of the five evaluation criteria in Table 1.
 They then rated each target text for each criterion on a five-point scale.
@@
 To collect overall judgments, we adopted a best-of-three approach, in which annotators are asked to choose the best text from three candidates.
-We did not use a five-point scale for the overall score because defining each rating level (from 1 to 5) for overall quality is difficult and could introduce bias Chiang et al.
+We did not use a five-point scale for the overall score because defining each rating level (from 1 to 5) for the overall quality is difficult and could introduce bias Chiang et al.
 (2024).
-Figure 4: Score distributions of the MMHE dataset.
+Figure 4: Human judgment score distributions for each task and criterion on the MMHE dataset.
 3.3 Data analysis
@@
 The correctness and completeness criteria exhibit diverse score distributions for most tasks, suggesting that even the state-of-the-art VLMs face challenges in these aspects.
-In contrast, the fluency criterion shows a narrow score distribution across most tasks, with a dominant score of 5.
+In contrast, the fluency criterion shows a narrow score distribution across most tasks, with a dominant score of five.
 This indicates that the VLMs generate fluent text even when visual understanding is inaccurate.
 Nonetheless, we consider fluency an essential criterion, as overlooking it could lead to high scores being assigned to correct but non-fluent text.
-A qualitative example illustrating this is provided in Section 4.3.
 The clarity and conciseness criteria have score distributions that are intermediate between the diverse and narrow distributions.
-While the score of 5 is more prevalent than in the diverse distributions, other scores still occur with noticeable frequency.
+While the score of five is more prevalent than in the diverse distributions, other scores still occur with noticeable frequency.
 Method REG VQA VDU IC Avg.
-Reference-based BLEU 45.3 29.4 57.3 46.8 44.7
+BLEU 45.3 29.4 57.3 46.8 44.7
 ROUGE 49.0 30.8 56.0 47.9 45.9
@@
 BART-S 56.4 20.5 60.9 57.8 48.9
-Ref.-free CLIP-S 60.1 39.7 60.9 52.0 53.2
+CLIP-S 60.1 39.7 60.9 52.0 53.2
 FLEUR 62.9 76.4 60.9 73.9 68.5
-HarmonicEval (ours) 66.6 76.4 73.4 77.0 73.4
-Table 2: Overall accuracy (%) on MMHE. The best result for each task is marked in bold. Average (Avg.) indicates the average performance across the four tasks.
-Metric REG VQA VDU IC
-Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc
+GPT-FLEUR 60.1 75.0 76.5 76.0 71.9
+HarmonicEval 66.6 76.4 73.4 77.0 73.4
+Table 2: Accuracy (%) on MMHE. The best result for each task is marked in bold. Average (Avg.) indicates the average accuracy across the four tasks.
+REG VQA VDU IC
+Metric Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc
 BLEU $6.0$ $6.9$ $3.9$ $1.2$ $6.1$ $1.3$ $10.4$ $11.0$ $19.3$ $4.1$ $19.8$ $12.9$ $14.9$ $14.3$ $21.2$ $4.4$ $4.5$ $5.9$ $0.3$ $11.3$
@@
 $24.5$ $14.0$
+GPT-FLEUR 19.7 $30.6$ 14.0 19.7 $11.0$ $\bm{54.7}$ $42.0$ $18.0$ $35.0$ $23.0$ 59.0 44.4 43.8 37.2 29.5
+$\bm{47.5}$
+47.7 25.3 29.1 13.4
 HarmonicEval
@@
 $\bm{23.8}$
-$\bm{53.5}$
+53.5
 $\bm{50.6}$
@@
 $\bm{45.8}$
-$\bm{44.7}$
+$44.7$
 $\bm{50.3}$
@@
 $\bm{22.8}$
-Table 3: Criterion-wise correlations on MMHE.
-Scores for the most positively and negatively correlated criteria are marked with black and gray underlines, respectively.
+Table 3: Criterion-wise correlation analysis on MMHE.
+Scores for the most positively and negatively correlated criteria are marked with red and blue underlines, respectively.
+The highest correlations for each criterion are highlighted in bold.
 Crt: Correctness, Cmp: Completeness, Clr: Clarity, Flu: Fluency, Cnc: Conciseness.
 4 Experiments
-4.1 Settings
-Baselines.
-We implemented eight baseline metrics for the experiments on MMHE.
-Four of these are traditional n-gram-based metrics: BLEU Papineni et al.
-(2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), and CIDEr Vedantam et al.
-(2015).
-The remaining three metrics leverage neural network-based approaches: BERTScore Zhang et al.
-(2020), BARTScore Yuan et al.
-(2021), and CLIPScore Hessel et al.
-(2021).
-Additionally, we employ FLEUR Lee et al.
-(2024b), a VLM-based evaluation metric that provides a single overall score and its textual explanation.
-Details of these baselines are provided in Appendix E.
-Performance measures.
-To assess the performance of evaluation metrics on MMHE, we use accuracy (%) for overall evaluations and Kendall’s tau correlation coefficient $\tau$ for criterion-wise evaluations.
-Implementation details.
-The GPT-4o model is used as a backbone of HarmonicEval.
-The temperature hyperparameter is set to zero so that the outputs are as deterministic as possible for reproducibility.
-More details are in Appendix C.
-4.2 Main results
-Overall performance.
-Table 2 compares HarmonicEval with eight metrics in terms of overall performance on MMHE.
-HarmonicEval achieves the highest accuracy on REG (66.6), VQA (76.4), VDU (73.4), and IC (77.0).
-While FLEUR also achieves the highest score of 76.4 on VQA, it falls short on VDU.
-These results underscore the effectiveness of HarmonicEval across multiple tasks.
-Criterion-wise performance.
-Table 3 shows the correlations between predicted scores and human judgments for each evaluation criterion.
-HarmonicEval evaluates numerical explainability by reporting correlations between its predicted criterion-wise scores and human judgment scores, whereas other metrics report correlations between their predicted overall scores and criterion-wise human judgment scores.
-As shown, HarmonicEval achieves the highest correlation across most criteria, highlighting its numerical explainability.
-FLEUR emerges as the second-best metric, a trend also observed in overall comparisons.
-Task-wise trends in baseline metrics.
-To gain deeper insights into the evaluation behavior of baseline metrics, we examine the task-wise correlations.
-Specifically, we explore which evaluation criteria baseline metrics tend to prioritize or deprioritize for each task, potentially hindering comprehensive evaluation.
-In Table 3, black underlines denote the most correlated criterion, while gray underlines indicate the least correlated criterion for each metric within a task.
+4.1 Performance on MMHE
+We evaluate the performance of HarmonicEval on MMHE and compare it with conventional metrics. We also examine how existing metrics prioritize or deprioritize specific criteria in each task.
+Settings.
+To assess the performance of each metric, we use accuracy (%) for the overall evaluation and the Kendall’s tau correlation coefficient $\tau$ for criterion-wise evaluations.
+We implement eight baselines, including four n-gram-based metrics Papineni et al.
+(2002); Lin (2004); Vedantam et al.
+(2015); Banerjee and Lavie (2005) and four neural network-based metrics Zhang et al.
+(2020); Yuan et al.
+(2021); Hessel et al.
+(2021); Lee et al.
+(2024b), as listed in two groups in Table 2.
+Among them, FLEUR Lee et al.
+(2024b) is a state-of-the-art VLM-based metric.
+Note that all baseline metrics don’t support criterion-wise scoring and produce only overall scores.
+The details of baselines are in Appendix E.
+HarmonicEval utilizes GPT-4o as its backbone.
+For a fair comparison with FLEUR, we also implement GPT-FLEUR, which substitutes GPT-4o for the original LLaVA-1.5-13B.
+Main results.
+Table 2 compares HarmonicEval with conventional metrics in terms of overall performance on MMHE.
+HarmonicEval achieves the highest accuracy in REG (66.6), VQA (76.4), IC (77.0), and attains the top average score of 73.4 across tasks.
+While GPT-FLEUR obtains the highest score on VDU (76.5), it performs less effectively on REG.
+These results underscore the strong multi-task capability of HarmonicEval.
+Correlation analysis.
+To investigate the behavior of each existing evaluation metric, we show the correlation between the predicted overall scores and the human judgment scores for each criterion in Table 3.
+This analysis reveals which evaluation criteria conventional metrics tend to prioritize or deprioritize for each task, potentially hindering a comprehensive evaluation.
+In the table, red and blue underlines denote the most and least correlated criterion, respectively, for each task.
+We observe task-wise trends.
 For REG, completeness shows the highest correlation across most metrics.
-This is understandable, as REG requires explicit expressions to identify a unique object by distinguishing it from surrounding objects.
+This is reasonable, as REG requires explicit expressions to identify a unique object by distinguishing it from surrounding objects.
 For VQA, most metrics are more strongly correlated with conciseness but less so with completeness.
-This indicates that existing metrics deprioritize completeness, potentially leading to inaccurate evaluations of insufficient answers.
-A similar trend is observed in VDU, where completeness is also deprioritized by existing metrics.
+This indicates that conventional metrics deprioritize completeness, potentially leading to inaccurate evaluations of insufficient answers.
+A similar trend is observed in VDU, where completeness is also deprioritized by conventional metrics.
 For IC, fluency exhibits low correlations for most metrics, suggesting a tendency to assign high scores even to texts lacking natural flow.
-A qualitative example illustrating this tendency is provided in Section 4.3.
-4.3 Analysis.
+Overall, these results underscore the necessity of a comprehensive evaluation metric in multi-task scenarios.
+Table 3 also shows the correlations between the criterion-wise scores of HarmonicEval and the human judgment scores.
+As shown, HarmonicEval achieves the highest correlation across most criteria, highlighting its ability for criterion-wise scoring.
+4.2 Analysis
 Figure 5: Qualitative examples.
-Qualitative examples.
-We present a qualitative comparison between HarmonicEval and FLEUR. To generate textual explanations for HarmonicEval, we prompt the VLM with “Why? Tell me the reason.”, following Lee et al. (2024b).
-Figure 5 illustrates two examples from the IC task.
-In Figure 5 (a), HarmonicEval detects severe fluency issues and reflects these in the overall score, whereas FLEUR fails to identify such deficiencies.
-This highlights that FLEUR may overlook specific evaluation criteria, such as fluency, as discussed in Section 1.
-In Figure 5 (b), HarmonicEval identifies errors in the caption more accurately than FLEUR, assigning a low correctness score and providing a clear, criterion-specific explanation.
-This suggests that HarmonicEval is able to offer not only numerical explainability but also textual explanation.
+Can HarmonicEval improve explainability?
+Providing feedback on evaluation results to users is important.
+To investigate whether HarmonicEval offers clear textual explanations, we prompt the VLM with “Why?
+Tell me the reason.” after calculating the overall scores, following Lee et al.
+(2024b).
+Figure 5 shows a qualitative comparison of HarmonicEval and FLEUR in the IC task.
+Panel (a) shows that HarmonicEval detects severe fluency issues and reflects this in its overall score, whereas FLEUR overlooks such deficiencies.
+This implies that FLEUR may miss certain criteria (e.g., fluency), as discussed in Section 3.3.
+Panel (b) demonstrates that HarmonicEval identifies incorrect details in the caption more precisely than FLEUR, assigning a low correctness score along with a criterion-specific explanation.
+These examples suggest that HarmonicEval can complement overall scores with informative textual feedback.
 Metric REG VQA VDU IC Avg.
-HarmonicEval $66.6$ $\bm{76.4}$ $73.4$ $\bm{77.0}$ $\bm{73.4}$
+HarmonicEval $\bm{66.6}$ $\bm{76.4}$ $73.4$ $\bm{77.0}$ $\bm{73.4}$
 w/o criterion-wise scoring
@@
 $65.7$ $75.0$ $73.4$ $76.5$ $72.6$
-w/o score smoothing
-$\bm{67.5}$ $70.5$ $70.4$ $72.4$ $70.2$
-Table 4: Ablation study. In the "w/o criterion-wise scoring" setting, we prompt the VLM to predict the overall score directly.
+Table 4: Ablation study.
 $\gamma$ REG VQA VDU IC Avg.
@@
 Hyperparameter study.
-VLM REG VQA VDU IC Avg.
-LLaVA-7B
-$62.9$ $72.0$ $61.4$ $73.4$ $67.4$
-LLaVA-13B
-$64.8$ $\bm{77.9}$ $63.2$ $72.9$ $69.7$
-GPT-4o
-$\bm{66.6}$ $76.4$ $\bm{73.4}$ $\bm{77.0}$ $\bm{73.4}$
-Table 6: Overall accuracy (%) for different VLMs.
-Ablation study.
-Table 4 highlights the effectiveness of three key components in HarmonicEval: criterion-wise scoring, score smoothing, and harmonic weighting.
-As shown, omitting criterion-wise scoring reduces performance on REG, VQA, and IC, indicating that comprehensive evaluation improves overall accuracy.
-Similarly, omitting harmonic weighting or score smoothing lowers performance across most tasks, demonstrating that these techniques stabilize the evaluation process.
-VLM REG VQA VDU IC
-Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc
-LLaVA-7B
-$14.5$ $14.0$ $14.8$ $\bm{28.4}$ $2.9$ $42.2$ $49.4$ $23.7$ $45.2$ $18.1$ $29.7$ $40.3$ $28.6$ $44.9$ $7.2$ $28.8$ $31.5$ $7.8$ $34.8$ $10.5$
-LLaVA-13B
-$\bm{31.4}$ $28.2$ $22.4$ $26.1$ $15.0$ $49.4$ $\bm{57.8}$ $26.7$ $49.7$ $30.5$ $39.7$ $45.9$ $36.4$ $42.2$ $27.7$ $35.6$ $40.0$ $4.5$ $27.4$ $5.2$
-GPT-4o $23.2$ $\bm{30.8}$ $\bm{24.0}$ $20.7$ $\bm{23.8}$ $\bm{53.5}$ $50.6$ $\bm{31.8}$ $\bm{51.9}$ $\bm{44.4}$ $\bm{60.0}$ $\bm{48.8}$ $\bm{47.9}$ $\bm{51.2}$ $\bm{45.8}$ $\bm{44.7}$ $\bm{50.3}$ $\bm{19.8}$ $\bm{36.4}$ $\bm{22.8}$
-Table 7: Criterion-wise performance for different VLMs.
-Hyperparameter study.
-Table 5 shows the hyperparameter study for $\gamma$.
-As expected, harmonic weighting using $1.00.5\leq\gamma<$ 1.0 performs better than uniform weighting ($\gamma=1.0$).
-Mathematical discussion underlying assumptions of these weightings is provided in Appendix A.
-Backbone VLMs.
-Tables 6 and 7 compare the overall performance and criterion-wise performance of three backbone VLMs: LLaVA-1.5-7B, LLaVA-1.5-13B, and GPT-4o.
-As shown, LLaVA-1.5-13B consistently perform better than LLaVA-1.5-13B.
-GPT-4o outperforms both models but leaves room for improvement, underscoring the need for future research to enhance alignment with human evaluations.
-Metric Flickr8k-EX Flickr8k-CF
-Reference-based BLEU 30.8 16.9
-ROUGE 32.3 19.9
-METEOR 41.8 22.2
-CIDEr 43.9 24.6
-SPICE 44.9 24.4
-BERT-S 39.2 22.8
-BERT-S++ 46.7
-TIGEr 49.3
-ViLBERTS-F 50.1
-FAIEr-4 52.6 35.4
-RefCLIP-Score 53.0 36.4
-Polos $\bm{56.4}$ 37.8
-RefFLEUR 51.9 38.8
-Reference-free UMIC 46.8
-FAIEr-r 50.1 32.4
-CLIP-S 51.5 34.4
-InfoCLIP 32.6 23.5
-InfoMetIC 54.2 36.3
-InfoMetIC+ 55.5 36.6
-FLEUR 53.0 38.6
-HarmonicEval $53.1$ $\bm{39.2}$
-Table 8: Comparison on existing human evaluation datasets for IC task.
-Performance on other IC datasets.
-To verify the generalizability of HarmonicEval’s effectiveness across different datasets, we additionally evaluated it on two human evaluation datasets: Flickr8k-EX and Flickr8k-CF.
-These datasets, widely used for image captioning, allow comparisons with a broader range of baselines, including learning-based metrics specifically designed for IC  [^2] We use Kendall’s tau-c and tau-b for Flickr8k-EX and Flickr8k-CF, respectively, following prior studies Hessel et al.
+VLM Method REG VQA VDU IC Avg.
+L-7B FLEUR $\bm{69.4}$ $69.1$ $\bm{63.2}$ $72.6$ $\bm{68.6}$
+Harmonic $62.9$ $\bm{72.0}$ $61.4$ $\bm{73.4}$ $67.4$
+L-13B FLEUR $62.9$ $76.4$ $60.9$ $\bm{73.9}$ $68.5$
+Harmonic $\bm{64.8}$ $\bm{77.9}$ $\bm{63.2}$ $72.9$ $\bm{69.7}$
+G-4o FLEUR $60.1$ $75.0$ $\bm{76.5}$ $76.0$ $71.9$
+Harmonic $\bm{66.6}$ $\bm{76.4}$ $73.4$ $\bm{77.0}$ $\bm{73.4}$
+Table 6: Comparison of HarmonicEval and FLEUR across different VLMs.
+Is each component in HarmonicEval essential?
+We conduct an ablation study for the two key components of HarmonicEval: criterion-wise scoring and harmonic weighting.
+Specifically, we examine two alternative approaches: 1) prompting the VLM to directly predict the overall score based on the five criteria, without using criterion-wise scores (w/o criterion-wise scoring); and 2) computing the overall score as a simple average of the criterion-wise scores (w/o harmonic weighing).
+Table 4 shows that both components improve overall performance.
+Removing criterion-wise scoring lowers scores on REG, VQA, and IC, indicating that explicit scoring for each criterion is vital for comprehensive evaluation.
+Similarly, omitting harmonic weighting reduces performance in most tasks, validating the effectiveness of our approach.
+Is statistical aggregation effective?
+Table 5 shows a hyperparameter study for $\gamma$. As expected, harmonic weighting with $\gamma=0.75$ performs the best. This justifies the importance of the statistical aggregation process.
+Is HamonicEval effective across various backbone VLMs?
+Table 6 shows the overall performance comparison between HarmonicEval and FLEUR using LLaVA-1.5-7B, LLaVA-1.5-13B, and GPT-4o as backbone models.
+HarmonicEval consistently outperforms FLEUR on both LLaVA-1.5-13B and GPT-4o.
+On the other hand, FLEUR slightly outperforms HarmonicEval when using LLaVA-1.5-7B.
+We found that this is because LLaVA-1.5-7B tends to underrate texts in the conciseness criterion compared to human-assigned scores.
+Nonetheless, these results highlight HarmonicEval’s effectiveness across different VLMs, particularly with more capable ones.
+Metric F-EX F-CF Com Pas
+Reference-based BLEU 30.8 16.9 30.6 72.9
+ROUGE 32.3 19.9 32.4 74.1
+METEOR 41.8 22.2 38.9 78.0
+CIDEr 43.9 24.6 37.7 76.8
+SPICE 44.9 24.4 40.3 69.6
+BERT-S 39.2 22.8 30.1 79.1
+TIGEr 49.3 45.4 80.7
+ViLBERTS-F 50.1 52.4 79.6
+FAIEr-4 52.6 35.4 57.7 81.4
+RefCLIP-Score 53.0 36.4 55.4 83.1
+Polos $\bm{56.4}$ 37.8 57.6 $\bm{86.5}$
+RefFLEUR 51.9 38.8 64.2 85.5
+Reference-free UMIC 46.8 30.1 56.1 85.1
+FAIEr-r 50.1 32.4 50.5
+CLIP-S 51.5 34.4 53.8 80.7
+InfoCLIP 32.6 23.5 15.3 64.1
+InfoMetIC 54.2 36.3 59.2 85.3
+InfoMetIC+ 55.5 36.6 59.3 $\bm{86.5}$
+FLEUR 53.0 38.6 63.5 83.2
+GPT-FLEUR 53.5 39.0 61.5 82.6
+HarmonicEval 53.1 $\bm{39.2}$ $\bm{66.2}$ 82.4
+Table 7: Performance comparison on the Flickr8k-Expert (F-EX), Flickr8k-CF (F-CF), Composite (Com) and Pascal-50S (Pas) datasets.
+Metric 1-ref 4-ref
+Polos 93.3 95.4
+RefFLEUR 97.3 $\bm{98.4}$
+FLEUR 96.8 96.8
+GPT-FLEUR 97.0 97.0
+HarmonicEval $\bm{97.8}$ 97.8
+Table 8: Performance comparison on the FOIL dataset.
+4.3 Performance on conventional IC datasets
+To assess the robustness of HarmonicEval in standard image captioning scenarios, we also evaluate it on five widely used IC datasets: Flickr8k-Expert, Flickr8k-CF, Composite, Pascal-50S, and FOIL.
+These datasets allow comparisons with a broader set of baselines, including specialized IC metrics222 Following prior works Hessel et al.
 (2021); Lee et al.
-(2024b) The details of additional baselines Anderson et al.
+(2024b), we use Kendall’s tau-c for Flickr8k-EX and Composite, tau-b for Flickr8k-CF, and accuracy for Pascal-50S.
+Additional baselines Anderson et al.
 (2016); Jiang et al.
@@
 (2023); Wada et al.
-(2024) are in Appendix E..
-As shown in Table 8, HarmonicEval achieves performance comparable to or even better than state-of-the-art metrics on the three datasets.
-While our main focus is on automatic multi-task, multi-criteria evaluation, these results highlight the robustness of HarmonicEval across multiple datasets.
+(2024) are described in Appendix E..
+As shown in Table 7 and 8, HarmonicEval achieves performance comparable to or even better than state-of-the-art metrics on all five datasets.
+Although our primary goal is a multi-task evaluation, these results highlight HarmonicEval’s robustness even in single-task settings.
 5 Related Work
@@
 Traditional metrics such as BLEU Papineni et al.
-(2002) and ROUGE Lin (2004) were originally proposed for automatic evaluation of machine translation and text summarization models.
+(2002) and ROUGE Lin (2004) were originally proposed for the automatic evaluation of machine translation and text summarization models.
 These metrics rely on n-gram overlap with reference texts.
@@
 (2023b) introduced a form-filling paradigm for criterion-based evaluation for text summarization and dialogue generation.
-Our HarmonicEval further extends this line of research with its capability of handling various vision-language tasks and provides a robust framework for aggregating criterion-wise scores based on the second-order statistics of token probability distributions assigned by VLMs.
-For vision-language tasks, CIDEr Vedantam et al.
+Our HarmonicEval further extends this line of research with its capability of handling various multi-modal tasks and provides a robust framework for aggregating criterion-wise scores based on the second-order statistics of token probability distributions assigned by VLMs.
+For the IC task, CIDEr Vedantam et al.
 (2015) measures the consensus between candidate and reference captions by weighting n-grams using TF-IDF.
@@
 (2024b), a state-of-the-art reference-free metric, provides textual explanations that underlie its overall scores.
-However, existing metrics primarily focus on providing overall evaluation, lacking numerical explainability and often overlooking specific criteria.
-HarmonicEval addresses these issues by offering criterion-wise scores alongside an overall score, enabling more interpretable and comprehensive evaluations.
+However, existing metrics primarily focus on providing the overall evaluation, and often overlook specific criteria.
+HarmonicEval addresses these issues by offering criterion-wise scores alongside the overall score, enabling comprehensive evaluation.
 Human evaluation datasets.
-Several human evaluation datasets have been developed for NLG tasks.
-For instance, the WMT shared tasks Semenov et al.
-(2023) provide benchmarks for machine translation, TMU-GFM Yoshimura et al.
-(2020) for grammatical error correction, and SummEval Fabbri et al.
-(2021) for text summarization.
-In the vision-language domain, existing datasets such as Flickr8k Hodosh et al.
+Several human evaluation datasets have been proposed for NLG tasks.
+Example datasets include SummEval Fabbri et al.
+(2021) for text summarization and the WMT shared tasks Semenov et al.
+(2023) for machine translation.
+In the multi-modal domain, existing datasets such as Flickr8k Hodosh et al.
 (2013), PASCAL-50S Vedantam et al.
@@
 (2024) target image captioning.
-In contrast, our MMHE dataset provides a multi-task, multi-criteria human evaluation resource, including sentences generated by several state-of-the-art VLMs across diverse vision-language tasks.
+In contrast, our MMHE dataset provides a multi-task, multi-criteria human evaluation resource, including sentences generated by several state-of-the-art VLMs across diverse tasks.
 This dataset enables more nuanced evaluations that capture a variety of criteria and tasks, advancing the field beyond the focus on captioning alone.
 6 Conclusion
-We introduced HarmonicEval and MMHE, a novel metric and dataset for automatic multi-criteria evaluation for various vision-language tasks, respectively.
-HarmonicEval is a reference-free evaluation metric that utilizes score variance to aggregate individual scores in a statistically reliable way.
-MMHE is a dataset covering four vision-language tasks and five criteria.
-Experimental results demonstrated the effectiveness of HarmonicEval on both MMHE and existing human evaluation datasets.
-We hope our study contributes to advancing research in automatic evaluation metrics.
+We introduced HarmonicEval, a novel reference-free evaluation metric for multiple multi-modal tasks.
+HarmonicEval predicts criterion-wise scores and aggregates them via a statistically principled method to produce an overall score.
+In addition, we constructed MMHE, the first multi-task multi-criteria human evaluation dataset, consisting of 18,000 expert judgments.
+Experimental results demonstrate that HarmonicEval aligns more closely with human judgments than existing metrics on both MMHE and other commonly used human evaluation datasets.
+Furthermore, our analysis with MMHE reveals that existing metrics tend to prioritize certain criteria while neglecting others.
 Limitations
@@
 Criteria.
-We carefully selected five general criteria that are considered effective across various vision-language tasks as detailed in Appendix B.
-These criteria were useful for discussions spanning the four vision-language tasks in this study.
+We carefully selected five general criteria that are considered effective across various multi-modal tasks as detailed in Appendix B.
+These criteria were useful for discussions spanning the four multi-modal tasks in this study.
 To further expand research to include a greater number of criteria and tasks, analyzing the relationship between task- or domain-specific criteria and these general criteria would also be necessary in future work.
@@
 Proposition 4. Harmonic weighting approaches to uniform weighting when $\gamma\to 1$.
-We believe Assumption 3 is reasonless, and this justifies why $1.00.5<\gamma<$ 1.0 is used in practice.
+We believe Assumption 3 is reasonable, and this justifies why $1.00.5<\gamma<$ 1.0 is used in practice.
 Appendix B Evaluation Criteria
-We selected five key criteria essential for evaluating the output text of vision-language tasks: correctness, completeness, clarity, fluency, and conciseness.
+We selected five key criteria essential for evaluating the output text of multi-modal tasks: correctness, completeness, clarity, fluency, and conciseness.
 Correctness and completeness assess the alignment between the task input (e.g., an image for image captioning (IC), or an image and question for visual question answering (VQA)) and the task output (e.g., a candidate caption for IC, or a candidate answer for VQA).
@@
 While coherence and grammar were considered, they were not included: coherence is more suitable for tasks involving long-form text, such as summarization, and grammar is more relevant for tasks emphasizing grammatical correctness, such as grammatical error correction.
-To validate these five criteria, we manually examined 100 instances across various vision-language tasks and confirmed their sufficiency for evaluating output text.
+To validate these five criteria, we manually examined 100 instances across various multi-modal tasks and confirmed their sufficiency for evaluating output text.
 Appendix C Implementation details
@@
 In Table 7, only the correctness and comprehensiveness prompts are used because datasets contain only short captions and for computational efficiency.
-In Tables 6 and 7, the prompts are slightly modified for each VLM to ensure that only the score is returned as the response.
+In Table 6, the prompts are slightly modified for each VLM to ensure that only the score is returned as the response.
 C.2 Prompts for target text generation
@@
 It automatically constructs training examples from captions to produce detailed token-level scores.
-Appendix F Computational Budget
-We run all the experiments on ABCI (https://abci.ai/), Compute Node(A), whose CPUs are two Intel Xeon Platinum 8360Y, and GPUs are eight NVIDIA A100 SXM4.
-The approximate total processing time is 40 hours.
+Appendix F Detailed experimental results
+Table 9 shows the full results of performance comparison on Flickr8k-Expert, Flickr8k-CF, Composite, and Pascal-50S.
+For Pascal-50S, we report category-wise correlations (HC, HI, HM, MM) and their average (Avg.).
+Metric Flickr8k-EX Flickr8k-CF Composite Pascal-50S
+HC HI HM MM Avg.
+Reference-based BLEU 30.8 16.9 30.6 53.0 92.4 86.7 59.4 72.9
+ROUGE 32.3 19.9 32.4 51.5 94.5 92.5 57.7 74.1
+METEOR 41.8 22.2 38.9 56.7 97.6 94.2 63.4 78.0
+CIDEr 43.9 24.6 37.7 53.0 98.0 91.5 64.5 76.8
+SPICE 44.9 24.4 40.3 52.6 93.9 83.6 48.1 69.6
+BERT-S 39.2 22.8 30.1 65.4 96.2 93.3 61.4 79.1
+TIGEr 49.3 45.4 56.0 99.8 92.8 74.2 80.7
+ViLBERTS-F 50.1 52.4 49.9 99.6 93.1 75.8 79.6
+FAIEr-4 52.6 35.4 57.7 59.7 $\bm{99.9}$ 92.7 73.4 81.4
+RefCLIP-Score 53.0 36.4 55.4 64.5 99.6 95.4 72.8 83.1
+Polos $\bm{56.4}$ 37.8 57.6 70.0 99.6 97.4 79.0 $\bm{86.5}$
+RefFLEUR 51.9 38.8 64.2 68.0 99.8 98.0 76.1 85.5
+Reference-free UMIC 46.8 30.1 56.1 66.1 99.8 $\bm{98.1}$ 76.2 85.1
+FAIEr-r 50.1 32.4 50.5
+CLIP-S 51.5 34.4 53.8 56.5 99.3 96.4 70.4 80.7
+InfoCLIP 32.6 23.5 15.3 37.3 87.3 58.9 72.9 64.1
+InfoMetIC 54.2 36.3 59.2 69.0 99.8 94.0 78.3 85.3
+InfoMetIC+ 55.5 36.6 59.3 $\bm{69.9}$ 99.7 96.8 $\bm{79.6}$ $\bm{86.5}$
+FLEUR 53.0 38.6 63.5 61.3 99.7 97.6 74.2 83.2
+GPT-FLEUR 53.5 39.0 61.5 59.4 99.5 97.3 74.3 82.6
+HarmonicEval 53.1 $\bm{39.2}$ $\bm{66.2}$ 58.7 99.8 97.0 74.2 82.4
+Table 9: Performance comparison on the Flickr8k-Expert, Flickr8k-CF, Composite, and Pascal-50S datasets.
+Appendix G Computational Budget
+We conduct all experiments on crowd compute nodes equipped with two Intel Xeon Platinum 8360Y CPUs and eight NVIDIA A100 SXM4 GPUs. The total processing time is approximately 40 hours.